### v2

# Redistribution Through

Abstract
We study how to optimally segment a monopolistic market given a redistributive objective. Optimal redistributive segmentations (i) induce the seller to price progressively, i.e., richer consumers pay higher prices than poorer ones, and (ii) may require giving a higher profit than uniform pricing if the redistributive motive is strong. We further show that optimal redistributive segmentations are implementable via price-based regulation.
JEL classification codes: D42, D83, D39.
Keywords: Market Segmentation; Price Discrimination; Redistribution; Information Design.
1 Introduction
Firms and online platforms use rich consumer data to segment markets and personalize prices. While these practices can improve economic efficiency and even increase total consumer surplus, questions remain about their distributional consequences. Since price discrimination affects consumers differently, assessing its welfare effects through aggregate consumer surplus conceals its heterogeneous impacts. In fact, segmenting markets so as to maximize total consumer surplus may disproportionately benefit wealthier consumers.
As Bergemann et al. (2015) show, monopolistic markets can be segmented so that the whole generated surplus accrues to consumers. This outcome is achieved by creating market segments that pool consumers with high and low willingness to pay. These segmentations exploit pooling externalities (Galperti et al., 2023): consumers with low willingness to pay exert a positive externality on consumers with high willingness to pay in the same segment by inducing the seller to charge prices below the uniform monopoly price. As a byproduct, buyers with a high willingness to pay capture most of the surplus. Nevertheless, a low willingness to pay might reflect adverse economic and social circumstances. Consumer surplus maximization may, therefore, unintentionally benefit wealthy buyers to the detriment of poorer ones.
In this paper, we study optimal segmentations for an objective that prioritizes the welfare of consumers with low willingness to pay. Our two main economic insights are the following. First, unlike segmentations that maximize total consumer surplus—which assign consumers with high willingness to pay to low-price segments—, optimal redistributive segmentations are positive-assortative: they assign buyers with high willingness to pay to high-price segments and buyers with low willingness to pay to low-price segments. As a result, they induce the monopolist to price progressively, meaning that wealthier consumers pay higher prices than poorer ones (Section 3.4). Second, this separation of wealthy and poor consumers into different segments might allow the seller to price discriminate more effectively and increase its profit compared to uniform pricing. Therefore, redistributive segmentations may inadvertently allocate part of the surplus created to the seller, thereby not maximizing total consumer surplus (Section 3.5).
Our model considers a continuous population of buyers, each having a unit demand for a good sold by a monopolistic seller. Each buyer is characterized by a type equal to their willingness to pay for the good. We refer to the initial distribution of types in the economy as the aggregate market. [^1] We follow the terminology of Bergemann et al. (2015). Without loss of generality, we assume that the seller’s cost of production is zero, so gains from trade are always positive. A market designer can flexibly segment the aggregate market, anticipating that the seller will charge a profit-maximizing price in each segment. The designer’s preferences are represented by a welfare function, which describes the social value of each buyer type consuming at each possible price. We restrict attention to welfare functions that align with consumer surplus: welfare is positive if and only if the buyer is willing to consume, and welfare is decreasing in price. Among those, we consider redistributive welfare functions: any price discount is more welfare-improving when granted to a low-type buyer rather than a high-type buyer. Mathematically, this property corresponds to supermodularity. Two examples of redistributive welfare functions are the weighted utilitarian welfare function with decreasing Pareto weights, which have received particular attention in the recent literature on redistributive mechanism design (Dworczak et al., 2021; Akbarpour ⓡ al., 2024b); and concave increasing transformations of consumer surplus, considered in the literature on optimal taxation (Mirrlees, 1971) and insurance (Rothschild and Stiglitz, 1976).
We begin by showing that optimal redistributive segmentations must be efficient (Lemma 1), so that they induce the seller to serve all buyers. Therefore, there is no equity-efficiency trade-off. [^2] This result is implied by the fact that any inefficient segmentation can be Pareto-improved (Haghpanah and Siegel, 2023). We then define the redistributive order, a partial order that ranks efficient segmentations according to their degree of redistributivity. Specifically, a segmentation is more redistributive than another if it yields higher aggregate welfare under any redistributive welfare function. Our first main result, Theorem 1, characterizes this order in terms of buyer mass transfers between market segments. We show that a segmentation is more redistributive than another if and only if it can be derived from the latter through a finite sequence of two kinds of transfers: (i) downward transfers, that move a certain mass of consumers to a lower-price segment; (ii) redistributive transfers, that swap a mass of buyers of a given type in a high-price segment with an equal mass of higher-type buyers in a lower-price segment. Downward transfers increase welfare because redistributive welfare functions decrease in price, whereas redistributive transfers increase welfare because redistributive welfare functions satisfy supermodularity. [^3] Our characterization of the redistributive order in terms of mass transfers is analogous to the characterization of the supermodular stochastic order carried out independently by Müller (2013) and Meyer and Strulovici (2015). For further discussion, see Section 3.2.
We then prove in Proposition 1 that the set of optimal redistributive segmentations is exactly the set of segmentations that are undominated in the redistributive order. Specifically, our definition of the redistributive order implies that, for any redistributive welfare function, an undominated optimal segmentation exists. Proposition 1 shows the converse statement: any undominated segmentation can be rationalized as optimal for some redistributive welfare function. Building on this equivalence, Proposition 2 further characterizes optimal redistributive segmentations. Theorem 1 implies that undominated segmentations are such that no downward or redistributive transfer can be performed without inducing the seller to charge a higher price in at least one segment. Proposition 2 identifies the set of prices between which the seller must be indifferent, so that performing any downward or redistributive transfer violates at least one of the seller’s incentive constraints.
We show the optimality of progressive pricing by proving that any optimal segmentation must assign buyer types to prices in a positive-assortative manner. Specifically, Theorem 2.A establishes that any optimal segmentation satisfies a weak form of monotonicity, which requires that, if a buyer consumes at a given price, there must exist a buyer with a greater willingness to pay who consumes at a higher price. Optimal segmentations may also satisfy a stronger notion of monotonicity: if a buyer consumes at a given price, then all buyers with a greater willingness to pay consume at a higher price. We show in Theorem 2.B that optimal segmentations satisfy this strong form of monotonicity if and only if the welfare function satisfies a strengthening of supermodularity reflecting strong redistributive concerns. Furthermore, optimal strongly monotone segmentations are obtained by following a simple greedy algorithm that iteratively maximizes the welfare of every type starting from the lowest.
Next, we show that redistributive segmentations may not maximize consumer surplus. The monotonicity of redistributive segmentations separates buyers with low and high willingness to pay into different segments, possibly allowing the seller to price-discriminate more effectively than under consumer-optimal segmentations. As a result, the seller might obtain a strictly positive share of the surplus generated by segmenting the market. In such cases, we say that the seller enjoys a redistributive rent. Redistributive rents arise especially when optimal segmentations satisfy the strong form of monotonicity described in the previous paragraph. Theorem 3 characterizes the markets for which strong redistributive preferences necessarily induce a redistributive rent for the seller. When rents are not necessary, the optimal segmentation is very simple with only two segments: one where the lowest possible price is charged, pooling all types that would not have bought the good in the unsegmented market; and one complementary segment where the uniform monopoly price is charged.
Our analysis considers the segmentation problem faced by a market designer with redistributive concerns. In practice, however, market segmentation is primarily conducted by private actors whose objectives are not necessarily aligned with welfare maximization. For instance, online platforms may focus on maximizing the number of transactions (Romanyuk and Smolin, 2019), while data intermediaries aim for profit maximization (Yang, 2022). In the last section of the paper, we hence investigate under what conditions redistributive segmentations can be implemented by a regulator who can only imperfectly observe the segmentation chosen by an agent, assumed here to be the seller. [^4] The result would still hold if the agent had a different objective function, provided it is aligned with efficiency. For example, a platform remunerated by a flat fee on transactions would seek to maximize participation and thus segment efficiently. We introduce a notion of price-based implementation: a segmentation is implementable through price-based regulation if there exists no alternative segmentation that strictly increases the seller’s profit while inducing the same distribution of prices. [^5] This notion of implementation is equivalent to Lin and Liu’s (2024) notion of credibility. We further discuss the connection to that contribution in Section 4, Footnote 28. This notion of implementation is motivated by the fact that although regulators are likely to possess less information than data intermediaries, they can plausibly observe the distribution of charged prices. [^6] Many websites, such as Camelcamelcamel, Keepa, Price Before, and Honey, already specialize in tracking the evolution, frequency and extent of price fluctuations for products sold on Amazon. In the context of regulating online platforms, for instance, it only involves monitoring transaction data and imposing penalties based on the divergence of the empirical price distribution from the desired target distribution. In Proposition 3, we show that any optimal redistributive segmentation is implementable through price-based regulation. [^7] In fact, Proposition 3 states a more general result of independent interest: any efficient segmentation can be implemented via price-based regulation.
Related literature
Our paper contributes to the analysis of the welfare effects of third-degree price discrimination, following Pigou (1920) and Robinson (1933). Much of this literature focuses on studying how given market segmentations affect efficiency, consumer surplus, total output and prices (Schmalensee, 1981; Varian, 1985; Aguirre et al., 2010; Cowan, 2016; Rhodes and Zhou, 2024). Le Grand (1975) is the first to study theoretically the use of third-degree price discrimination as a redistributive device for public firms.
Bergemann et al. (2015) pioneered an information-design-based approach to third-degree price discrimination by characterizing the feasible welfare outcomes under any segmentation of a monopolistic market. [^8] We refer to Kamenica (2019) and Bergemann and Morris (2019) for comprehensive reviews of the Bayesian persuasion and information design literature. They show that any combination of consumer surplus and producer profit such that the seller’s profit is at least equal to that under uniform pricing, the consumer surplus is non-negative, and the total surplus does not exceed the efficient level, is achievable through some segmentation. Haghpanah and Siegel (2022) and Terstiege and Vigier (2023) perform the same exercise in a multi-product setting, while Elliott et al. (2024) and Bergemann et al. (2024) consider competitive markets. Kartik and Zhong (2023), extend the characterization of welfare outcomes of Bergemann et al. to the case where the buyer and the seller have correlated values. [^9] A parallel literature following Roesler and Szentes (2017) assumes that buyers are imperfectly informed about their valuation and characterize the consumer-optimal information structures. Condorelli and Szentes (2020) characterize consumer-optimal information structures in the hold-up problem. Ravid et al. (2022) consider the case where the buyer has a cost of acquiring information. Deb and Roesler (2021) characterize consumer-optimal information structures in the multidimensional screening problem. Armstrong and Zhou (2022) characterize consumer-optimal information structures when the downstream market is a differentiated product duopoly. We contribute to this literature by studying the impact of market segmentation on consumer welfare beyond the standard of consumer surplus. In particular, we study redistributive welfare functions and show that prioritizing low-type consumers does not hinder efficiency but may not maximize total consumer surplus.
Doval and Smolin (2024) study the welfare effects of information provision on a heterogeneous population of agents. They introduce the Bayes welfare set, consisting of the set of welfare profiles that can be achieved under any information policy. They characterize this set geometrically and show that its Pareto frontier is generated by the set of solutions of a family of Bayesian persuasion problems, in which a weighted utilitarian social planner plays the role of the designer. As such, they characterize the welfare effects that any segmentation can have on consumers of different types, without assuming any particular shape for the welfare function. In our paper, we assume the welfare function satisfies particular conditions—specifically, redistributivity—, solve explicitly for the optimal segmentations and show that their structure differs from Bergemann et al.’s consumer-surplus maximizing segmentations.
Dubé and Misra (2023) conduct a randomized controlled field experiment to study the welfare implications of personalized pricing. Although they find a negative effect of personalized pricing on total consumer surplus, the effect on aggregate consumer welfare is positive under some inequality-averse welfare functions. This shows that low-valuation buyers are the ones who benefit from the implemented price discrimination. Buchholz et al. (2020) study the effect of price discrimination on a two-sided platform for cab rides. They also find that while aggregate consumer surplus decreases, more price-sensitive buyers benefit from personalized pricing.
Following the seminal contribution of Weitzman (1977), a burgeoning literature has integrated redistributive concerns into mechanism design (see, e.g., Condorelli, 2013; Reuter and Groh, 2020; Dworczak et al., 2021; Pai and Strack, 2022; Akbarpour ⓡ al., 2024a, b; Kang, 2023; Kang and Watt, 2024). [^10] Dworczak (2024) gives an overview of the recent contributions to inequality-aware mechanism design. Our paper studies redistribution in an information design setting. [^11] Arya and Malhotra (2022) studies redistribution through a different notion of market segmentation, namely, the non-tradability of essential and non-essential goods in a general equilibrium framework. Contrary to the optimal mechanisms in this literature, our redistributive segmentations are socially efficient and thus do not exhibit any equity-efficiency trade-off.
2 Model
Economy
A monopolistic seller (she) supplies a good to a unit mass of buyers. Buyers have a unit demand and differ in their willingness to pay for the good, $\theta$, assumed to belong to the finite set $\Theta=\{\theta_{1},\dots,\theta_{K}\}\subset\mathbb{R}_{+}$. We refer to the willingness to pay as the buyer’s type and use the two terms interchangeably. A buyer willing to pay $\theta$ consumes at a price $p\in\mathbb{R}_{+}$ if and only if $\theta\geq p$. We assume that the seller’s marginal cost of production is zero, so gains from trade are positive. [^12] This assumption is without loss of generality. If the seller’s marginal cost were equal to $c>0$, then one could redefine the buyer type as $\tilde{\theta}=\max\{0,\theta-c\}$ for all $\theta\in\Theta$. The aggregate market (in short the market) $\mu\in\Delta(\Theta)$ specifies the initial distribution of types.
Market segmentation
A market segmentation is a division of the population of buyers into different market segments indexed by $s\in S$, where $S$ is a sufficiently large finite set. Formally, a segmentation corresponds to a probability distribution $\sigma\in\Delta(\Theta\times S)$ describing the mass $\sigma(\theta,s)$ of type $\theta$ buyers assigned to segment $s$, and satisfying the following market consistency constraint:
$\forall\theta\in\Theta,\quad\sum_{s\in S}\sigma(\theta,s)=\mu(\theta).$ ($\mathrm{C}_{\mu}$)
Condition $\mathrm{C}_{\mu}$ ensures that the mass of type $\theta$ buyers assigned to all segments coincides with the mass of those types in the market. [^13] We can equivalently think of any pair $(\sigma,S)$ as a Blackwell experiment which, for any prior distribution $\mu$ over the state space $\Theta$, describes the joint distribution of signals $s\in S$ and states of the world $\theta\in\Theta$, and satisfies the Bayes-plausibility constraint $\mathrm{C}_{\mu}$ (Aumann et al., 1995; Kamenica and Gentzkow, 2011). According to that interpretation, each $\sigma$ represents a way of revealing information about the buyers’ types to the seller.
Seller’s pricing
When facing a market segmentation, the seller sets prices so as to maximize his profit on each generated segment. That is, for any $\mu\in\Delta(\Theta)$, any $\sigma\in\Sigma(\mu)$, and any $s\in S$ such that $\sigma(\theta,s)>0$ for some $\theta\in\Theta$, the seller solves
$\max_{p\in P}\;p\sum_{\theta\geq p}\sigma(\theta,s),$
where the set of prices is given by $P=\Theta$. Indeed, there is no loss of generality in assuming that the seller charges prices in $\Theta$ since, for any $k\in\{1,\dots,K\}$, charging a price falling strictly between two consecutive types $\theta_{k}$ and $\theta_{k+1}$ always yields a lower profit to the seller than charging either $p=\theta_{k}$ or [^14] It is also easy to see that charging any price lower than $\theta_{1}$ or greater than $\theta_{K}$ is strictly dominated. $p=\theta_{k+1}$.
Direct segmentations
Following Proposition 2 in Bergemann et al. (2015), it is without loss of generality to focus on direct market segmentations. A segmentation is direct if each price is charged in at most one segment. Direct segmentations correspond to direct communication mechanisms in the sense of Myerson (1982). In such mechanisms $S=P$, so that messages correspond to price recommendations to the seller.
A direct segmentation is obedient if the seller finds it optimal to follow the price recommendation in each segment. That is:
$\forall p,q\in P,\quad p\sum_{\theta\geq p}\sigma(\theta,p)\geq q\sum_{\theta\geq q}\sigma(\theta,p).$ (Ob)
For any $\mu\in\Delta(\Theta)$, we denote the set of direct and obedient segmentations of $\mu$ as $\Sigma(\mu)$. We henceforth slightly abuse terminology by referring to any element of $\Sigma(\mu)$ as a segmentation of $\mu$. For any $\mu\in\Delta(\Theta)$ and any $\sigma\in\Sigma(\mu)$, we let $\sigma_{P}\in\Delta(P)$ be the price distribution induced by $\sigma$, defined by:
$\forall p\in P,\quad\sigma_{P}(p)=\sum_{\theta\in\Theta}\sigma(\theta,p).$
For any $\mu\in\Delta(\Theta)$, we let $p^{\star}_{\mu}\in P$ be the uniform optimal price, that is, the price the seller would optimally charge in the aggregate market: [^15] Whenever there are several uniform optimal prices for some market $\mu\in\Delta(\Theta)$, we let $p^{\star}_{\mu}$ be the lowest one. This tie-breaking rule is inconsequential for any of our results.
$\forall p\in\operatorname{supp}(\mu),\quad p^{\star}_{\mu}\sum_{\theta\geq p^{\star}_{\mu}}\mu(\theta)\geq p\sum_{\theta\geq p}\mu(\theta).$
Let $\mu\in\Delta(\Theta)$. An instance of direct segmentation of $\mu$ is illustrated on Figure 1.
$\theta$$p$$(\theta_{1},\theta_{1})$$p^{\star}_{\mu}$$\theta_{n}$$p^{\star}_{\mu}$$\theta_{n}$$p^{\prime}$$p^{\prime\prime}$
Figure 1: A direct segmentation with two segments.
The initial mass of consumers, which is located at the row $p^{\star}_{\mu}$ in the grid $\Theta\times P$, is split between two segments $p^{\prime}$ and $p^{\prime\prime}$ in a vertical way, so that $\mathrm{C}_{\mu}$ is satisfied. Furthermore, Ob requires that the mass moved at each point must be such that $p^{\prime}$ and $p^{\prime\prime}$ are respectively profit-maximizing prices in segments $p^{\prime}$ and $p^{\prime\prime}$.
Redistributive welfare functions
We consider a designer (he) who can segment the market flexibly. His preferences over market outcomes are represented by a welfare function $w\colon\Theta\times P\to\mathbb{R}$. For any welfare function $w$ and segmentation $\sigma\in\Sigma(\mu)$, the aggregate welfare is
$\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma(\theta,p).$
The value $w(\theta,p)$ thus represents the individual contribution to the aggregate welfare when a buyer of type $\theta$ is assigned to a segment where the seller charges price $p$.
We restrict attention to welfare functions satisfying three conditions. First, the value from assigning a buyer to a segment where he does not consume is zero, and the value from assigning a buyer to a segment where he consumes is non-negative. That is, $w(\theta,p)=0$ if $\theta<p$, and $w(\theta,p)\geq 0$ if $\theta\geq p$. Second, assigning any buyer to a segment where the seller charges a lower price is always welfare-improving. That is, $p\mapsto w(\theta_{k},p)$ is a non-increasing function on $\{\theta_{1},\dots,\theta_{k}\}$ for all $k\in\{1,\dots,K\}$. The two preceding conditions capture the fact that social welfare aligns with consumer surplus. [^16] It is straightforward to verify that the buyer-utilitarian welfare function, given by $w(\theta,p)=(\theta-p)\mathds{1}_{\theta\geq p}$ for all $(\theta,p)\in\Theta\times P$, satisfies these two minimal conditions. The third condition captures the designer’s redistributive concerns. We require that any welfare function is supermodular conditional on buyers consuming. That is, for any $(\theta,p),(\theta^{\prime},p^{\prime})\in\Theta\times P$ such that $\theta^{\prime}>\theta\geq p^{\prime}>p$,
$w(\theta,p)-w(\theta,p^{\prime})\geq w(\theta^{\prime},p)-w(\theta^{\prime},p^{\prime}).$ (R)
Condition R ensures that any price discount increases welfare more when it is granted to low-type buyers rather than high-types.
We say that a welfare function is redistributive if it satisfies the three conditions above and denote as $\mathcal{R}$ the class of those welfare functions. Furthermore, we say that $w$ is strictly redistributive if it satisfies condition R strictly on all $\Theta\times P$ and is such that $p\mapsto w(\theta_{k},p)$ is strictly decreasing on $\{\theta_{1},\dots,\theta_{k}\}$ for all $k\in\{1,\dots,K\}$. We denote the class of strictly redistributive welfare functions as $\overline{\mathcal{R}}$.
In Examples 1 and 2 below, we highlight that the class of redistributive welfare functions encompasses two common specifications.
Example 1 (Decreasing Pareto weights).
Let $\lambda\colon\Theta\to\mathbb{R}$ be such that $\lambda(\theta)\geq 0$ for any $\theta\in\Theta$. The welfare function defined by
$\forall(\theta,p)\in\Theta\times P,\quad w(\theta,p)=\lambda(\theta)(\theta-p)\mathds{1}_{\theta\geq p},$ (1)
belongs to $\mathcal{R}$ whenever $\theta\mapsto\lambda(\theta)$ is non-increasing, and belongs to $\overline{\mathcal{R}}$ whenever $\theta\mapsto\lambda(\theta)$ is strictly decreasing. The class of welfare functions given by Equation 1 for some decreasing function $\lambda$ correspond to the family of weighted buyer-utilitarian welfare functions with decreasing Pareto weights. Under such welfare functions, the value to the designer of an additional unit of surplus is lower for high-types than for low-types. Decreasing Pareto weights can endogenously arise under a utilitarian objective if buyers with a low willingness to pay have a higher marginal utility of money than those with a high willingness to pay (Condorelli, 2013; Dworczak et al., 2021), or if buyers are identified by unobserved welfare weights that negatively correlate with their willingness to pay (Akbarpour ⓡ al., 2024b).
Example 2 (Increasing concave transformation).
Let $u\colon\mathbb{R}\to\mathbb{R}$ be an increasing function such that $u(0)=0$. The welfare function defined by
$\forall(\theta,p)\in\Theta\times P,\quad w(\theta,p)=u\bigl{(}(\theta-p)\mathds{1}_{\theta\geq p}\bigr{)},$ (2)
belongs to $\mathcal{R}$ if it is concave, and belongs to $\overline{\mathcal{R}}$ if it is strictly concave. Increasing concave transformations of surplus correspond to the class of welfare functions usually considered in optimal taxation (Mirrlees, 1971) and insurance (Rothschild and Stiglitz, 1976). [^17] Garcia and Tsur (2021) characterize optimal segmentations of a competitive insurance market. Such increasing concave transformations of surplus thus arise endogenously under a utilitarian objective. Concavity reflects equity concerns, as it implies diminishing marginal returns to welfare from additional units of surplus. Any designer whose welfare function is given by Equation 2 for some increasing concave $u$ would therefore prefer high prices to be charged to high-types than to low-types.
Note also that any welfare function which is a combination of Examples 1 and 2, given by
$\forall(\theta,p)\in\Theta\times P,\quad w(\theta,p)=\lambda(\theta)u\bigl{(}(\theta-p)\mathds{1}_{\theta\geq p}\bigr{)},$
belongs to $\mathcal{R}$ whenever $\theta\mapsto\lambda(\theta)$ is sufficiently decreasing, or $u$ is sufficiently concave, or both.
Segmentation design
For any given welfare function $w\in\mathcal{R}$ and an aggregate market $\mu\in\Delta(\Theta)$, the designer’s optimization program consists of choosing a segmentation of $\mu$ to maximize aggregate welfare:
$\max_{\sigma\in\Sigma(\mu)}\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma(\theta,p).$ ($\mathrm{P}_{w,\mu}$)
For any $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$, we say that $\sigma$ is $\mathrm{P}_{w,\mu}$-optimal if it solves the optimization program $\mathrm{P}_{w,\mu}$. Furthermore, we say that a segmentation is redistributive if there exist $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$ such that $\sigma$ is $\mathrm{P}_{w,\mu}$-optimal.
3 Redistributive segmentations and welfare
3.1 Efficiency
We start by proving that it is without loss of generality to search for optimal segmentations among the efficient ones. A segmentation is efficient if it leads the seller to serve all the buyers. We define $\Omega=\bigl{\{}(\theta,p)\in\Theta\times P\ |\ \theta\geq p\bigr{\}}$ as the set of market outcomes where buyers consume. Hence, a segmentation is efficient if it is supported on $\Omega$. For any $\mu\in\Delta(\Theta)$, we let
$\Sigma_{\Omega}(\mu)\coloneqq\bigl{\{}\sigma\in\Sigma(\mu)\ |\ \operatorname{supp}(\sigma)\subseteq\Omega\bigr{\}},$
be the set of efficient segmentations of $\mu$.
Lemma 1 (Efficiency).
For any $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$, there exists a $\mathrm{P}_{w,\mu}$-optimal segmentation that belongs to $\Sigma_{\Omega}(\mu)$.
Proof.
Let $\mu\in\Delta(\Theta)$ and consider $\sigma\in\Sigma(\mu)$ such that $\operatorname{supp}(\sigma)\not\subset\Omega$. There must therefore exist $(\theta,p)\in\operatorname{supp}(\sigma)$ such that $\theta<p$. Let us contemplate transferring the entire mass of buyers at $(\theta,p)$ to $(\theta,\theta)$. This transfer violates neither $\mathrm{C}_{\mu}$ nor Ob. Indeed, the total mass of type $\theta$ buyers still equals $\mu(\theta)$ so that condition $\mathrm{C}_{\mu}$ is preserved. Furthermore, this buyer transfer relaxes the obedience constraints in both segments $\theta$ and $p$. Finally, since $w(\theta,p)=0\leq w(\theta,\theta)$, the transfer must increase the aggregate welfare. In fact, aggregate welfare weakly increases without hurting the seller or the buyers—a (weak) Pareto improvement. ∎
Lemma 1 establishes that segmenting a market in a redistributive way never hinders efficiency. The intuition behind that result is similar to Haghpanah and Siegel’s (2023) main result. Namely, any inefficient segmentation is Pareto-dominated by some efficient one, which improves aggregate welfare as redistributive welfare functions are aligned with consumer surplus.
Operationally, Lemma 1 implies that there is no loss of generality in considering only efficient segmentations as candidates for $\mathrm{P}_{w,\mu}$-optimality. We can therefore restate the segmentation design problem $\mathrm{P}_{w,\mu}$ as choosing a welfare-maximizing efficient segmentation:
$\max_{\sigma\in\Sigma_{\Omega}(\mu)}\sum_{(\theta,p)\in\Omega}w(\theta,p)\,\sigma(\theta,p)$ ($\mathrm{P}^{\star}_{w,\mu}$)
3.2 The redistributive order
We propose an ordering of efficient segmentations based on their redistributivity, which we call the redistributive order. This ordering will help characterize qualitative properties of redistributive segmentations in Sections 3.3 and 3.4.
Definition 1.
Let $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$. We say that $\sigma$ is more redistributive than $\sigma^{\prime}$, denoted $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$, if
$\forall w\in\mathcal{R},\quad\sum_{(\theta,p)\in\Omega}w(\theta,p)\,\sigma(\theta,p)\geq\sum_{(\theta,p)\in\Omega}w(\theta,p)\,\sigma^{\prime}(\theta,p),$
We call $\succsim_{\mathcal{R}}$ the redistributive order over $\Sigma_{\Omega}(\mu)$ and denote as $\succ_{\mathcal{R}}$ its asymmetric part. [^18] One can show that $\succsim_{\mathcal{R}}$ is a partial order, that is, a transitive, reflexive and antisymmetric binary relation. While the first two conditions are easy to verify from the definition, the fact that it is antisymmetric follows from Theorem 1 below.
Two efficient segmentations are ranked according to the redistributive order if one leads to greater aggregate welfare for any redistributive welfare function. Given any market $\mu$, a more redistributive segmentation of $\mu$ is thus precisely one that is preferred ex-ante by any designer with a redistributive objective. Note that $\succsim_{\mathcal{R}}$ is not a complete order. One specific segmentation may produce higher aggregate welfare than another for certain redistributive welfare functions, while the latter produces higher aggregate welfare than the former for different redistributive welfare functions.
The redistributive order is defined in the same spirit as the (bivariate) supermodular stochastic order (Shaked and Shanthikumar, 2007, Section 9.A.4.), with two notable differences. First, in addition to being supermodular, redistributive welfare functions are also decreasing in their second dimension (corresponding here to the price/segment). Second, efficient segmentations correspond to bivariate distributions such that only the first marginal is prescribed. We leverage this connection by characterizing the redistributive order in a similar way Müller (2013) and Meyer and Strulovici (2015) characterize the supermodular stochastic order. Namely, we characterize $\succsim_{\mathcal{R}}$ in terms of mass transfers.
A buyer mass transfer (in short a transfer) is a function $t\colon\Omega\to\mathbb{R}$ describing buyers’ mass movements within $\Omega$. Among all the possible transfers, the following two classes will enable us to characterize $\succsim_{\mathcal{R}}$.
Definition 2 (Downward buyer transfer).
A buyer transfer $t\colon\Omega\to\mathbb{R}$ is downward if there exists $\delta\geq 0$ and $\tilde{\theta},p^{\prime},p^{\prime\prime}\in\Theta$ such that $p^{\prime}<p^{\prime\prime}$, and
$t(\theta,p)=\left\{\begin{array}[]{ll}\delta&\text{if $(\theta,p)=(\tilde{\theta},p^{\prime})$}\\ -\delta&\text{if $(\theta,p)=(\tilde{\theta},p^{\prime\prime})$}\\ 0&\text{otherwise}\end{array}\right.$
Definition 3 (Redistributive buyer transfer).
A buyer transfer $t\colon\Omega\to\mathbb{R}$ is redistributive if there exists $\varepsilon\geq 0$ and $(\theta^{\prime},p^{\prime\prime}),(\theta^{\prime\prime},p^{\prime})\in\Omega$ such that $\theta^{\prime}<\theta^{\prime\prime}$, $p^{\prime}<p^{\prime\prime}$, and
$t(\theta,p)=\left\{\begin{array}[]{ll}\varepsilon&\text{if $(\theta,p)\in\bigl{\{}(\theta^{\prime},p^{\prime}),(\theta^{\prime\prime},p^{\prime\prime})\bigr{\}}$}\\[5.0pt] -\varepsilon&\text{if $(\theta,p)\in\bigl{\{}(\theta^{\prime},p^{\prime\prime}),(\theta^{\prime\prime},p^{\prime})\bigr{\}}$}\\[5.0pt] 0&\text{otherwise}\end{array}\right.$
Figure 2 illustrates the downward and redistributive transfer operations on an arbitrary segmentation.
$\theta$$p$$\delta$$\varepsilon$$\varepsilon$
Figure 2: The shaded region represents the set $\Omega$. A downward transfer involving a mass $\delta$ of buyers (in blue) and a redistributive transfer involving a mass $\varepsilon$ of buyers (in orange).
Downward transfers carry a certain mass of a given type of consumer from one segment to another with a lower price. They are always desirable since the redistributive welfare functions decrease in price. Redistributive transfers exchange a mass of some type assigned to a particular segment for the same mass of higher types assigned to a segment with a lower price. These transfers are also always desirable since condition R ensures that the welfare gain from lowering prices for low-types outweighs the welfare loss from increasing prices for high-types.
Therefore, any combination of downward and redistributive transfers increases aggregate welfare if the welfare function is redistributive. As our first result shows, these combinations of transfers are also the only ones to increase aggregate welfare for any redistributive objective. Let $\mathcal{T}$ be the set of all downward and redistributive buyer transfers.
Theorem 1 (Redistributive order).
For any $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$, $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$ if and only if there exists $n\in\mathbb{N}$, $n\neq 0$, and $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that
$\sigma=\sigma^{\prime}+\sum_{i=1}^{n}t_{i}.$
Theorem 1 states that one segmentation is less redistributive than another if and only if one can move from the former to the latter by performing a finite number of the elementary transfers described in Definitions 2 and 3. Its proof, which can be found in Appendix A, relies on duality arguments and follows a similar approach to the proofs of Theorem 2.4.1 in Müller (2013) and Theorem 1 in Meyer and Strulovici (2015).
3.3 Order-maximal segmentations and optimality
We now leverage Theorem 1 to uncover properties of redistributive segmentations. For any $\mu\in\Delta(\Theta)$, we say that a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is $\succsim_{\mathcal{R}}$-maximal if there exists no other segmentation $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ such that $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$.
Proposition 1 (Optimality).
For any $\mu\in\Delta(\Theta)$, the following two statements are true.
  1. (i)
    For any $w\in\mathcal{R}$, there exists a $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation of $\mu$ that is $\succsim_{\mathcal{R}}$-maximal. Furthermore, if $w\in\overline{\mathcal{R}}$, then any $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation of $\mu$ is $\succsim_{\mathcal{R}}$-maximal.
  2. (ii)
    Conversely, if $\sigma\in\Sigma_{\Omega}(\mu)$ is $\succsim_{\mathcal{R}}$-maximal, then there exists $w\in\mathcal{R}$ such that $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal.
Proposition 1 establishes a correspondence between optimal segmentations and the maximal elements of the redistributive order. Claim (i) asserts that it is without loss of generality to restrict attention to $\succsim_{\mathcal{R}}$-maximal elements when looking for redistributive segmentations (and even necessary to do so when the welfare function is strictly redistributive). This implication is a direct consequence of Theorem 1, since for any non-maximal segmentation there exists at least one sequence of transfers in $\mathcal{T}$ which increases aggregate welfare for any redistributive welfare function without violating any of the constraints defining $\Sigma_{\Omega}(\mu)$. Claim (ii) states that for any aggregate market, a $\succsim_{\mathcal{R}}$-maximal element can be rationalized as an optimal segmentation for some redistributive welfare function. Its proof can be found in Appendix B and relies on a separation argument.
Building on Proposition 1, we show that the saturation of a particular subset of the seller’s obedience constraints characterizes redistributive segmentations. [^19] Haghpanah and Siegel (2024) use a similar notion to our downward transfers to characterize when a segmentation is stable (meaning that no coalition of buyers can obtain a better outcome by deviating to another segment). In particular, they define a segmentation as saturated if transferring any positive mass of buyers to a given segment from a segment with a higher price increases the price in the lower segment. They show that a segmentation is stable if and only if it is efficient and saturated.
Proposition 2 (Characterization).
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is $\succsim_{\mathcal{R}}$-maximal if and only if the following conditions hold:
  1. (i)
    At any segment $p\in\operatorname{supp}(\sigma_{P})$ such that $p<\max\operatorname{supp}(\sigma_{P})$, the seller is indifferent between charging $p$ and some other price $q>p$;
  2. (ii)
    For any segment $p\in\operatorname{supp}(\sigma_{P})$ and any $\theta\geq p$, if $\theta\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}$ then, at any segment $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}\leq\theta$, the seller is indifferent between charging $p^{\prime}$ and $\theta$.
The proof of Proposition 2 can be found in Appendix C. Let us explain its logic. From Theorem 1 we know that, for any $\mu\in\Delta(\Theta)$, a $\succsim_{\mathcal{R}}$-maximal segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is precisely one on which performing any sequence $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ of downward or redistributive transfers would imply that $(\sigma+\sum_{i=1}^{n}t_{i}\bigl{)}$ violates Ob. [^20] We can restrict attention to how downward and redistributive transfers impact the seller’s obedience constraints since, for any market $\mu\in\Delta(\Theta)$, segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ and transfer $t\in\mathcal{T}$ sufficiently small, the resulting distribution $\sigma+t$ must satisfy $\mathrm{C}_{\mu}$. Indeed, the transfers belonging to $\mathcal{T}$ preserve the total mass of each buyer type, as $\sum_{p\in P}t(\theta,p)=0$ for any $\theta\in\Theta$. We start by establishing in Lemma 6 that, for a segmentation to be $\succsim_{\mathcal{R}}$-maximal it is sufficient that any single transfer $t\in\mathcal{T}$ induces $\sigma+t$ to violate Ob. For any $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$, we thus say that a transfer $t$ is feasible for $\sigma$ if it preserves Ob, that is, if $(\sigma+t)\in\Sigma_{\Omega}(\mu)$. We conversely say that a transfer $t\in\mathcal{T}$ is unfeasible for $\sigma$ if it is not feasible.
Let us investigate the conditions under which a particular downward transfer is unfeasible for some segmentation. Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$ be such that there exist two segments $p^{\prime},p^{\prime\prime}\in\operatorname{supp}(\sigma_{P})$ with $p^{\prime\prime}>p^{\prime}$, and let $\tilde{\theta}\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p^{\prime\prime})\bigr{)}$. Consider transferring downward a mass $\delta$ of types $\tilde{\theta}$ from the segment $p^{\prime\prime}$ to $p^{\prime}$, as illustrated in Figure 3(a).
[…]
(a) Downward transfer.
[…]
(b) Redistributive transfer.
Figure 3: Points within the orange boxes represent $(\theta,p)$ for which Ob tightens after a transfer.
In segment $p^{\prime}$, including a mass $\delta$ of types $\tilde{\theta}$ increases the demand at any price $q\in\{p^{\prime},\dots,\tilde{\theta}\}$ by an amount $\delta$ while keeping the demand at any other price unaffected. Therefore, this transfer leads the seller to deviate in segment $p^{\prime}$ to some price $q$ with $p^{\prime}<q\leq\tilde{\theta}$ if and only if
$\delta>\frac{1}{q-p^{\prime}}\Biggl{(}p^{\prime}\sum_{\theta\geq p^{\prime}}\sigma(\theta,p^{\prime})-q\sum_{\theta\geq q}\sigma(\theta,p^{\prime})\Biggr{)}.$ (3)
In segment $p^{\prime\prime}$, excluding a mass $\delta$ of types $\tilde{\theta}$ decreases the demand at any price $q\in\{p^{\prime},\dots,\tilde{\theta}\}$ by an amount $\delta$. Therefore, the seller deems it profitable to deviate in segment $p^{\prime\prime}$ to some price $q$ with $q>\tilde{\theta}$ if and only if
$\delta>\frac{1}{p^{\prime\prime}}\Biggl{(}p^{\prime\prime}\sum_{\theta\geq p^{\prime\prime}}\sigma(\theta,p^{\prime\prime})-q\sum_{\theta\geq q}\sigma(\theta,p^{\prime\prime})\Biggr{)}.$ (4)
It is clear that no downward transfer of type $\tilde{\theta}$ from segment $p^{\prime\prime}$ to segment $p^{\prime}$ is feasible for $\sigma$ if, in one of the two segments $p^{\prime}$ and $p^{\prime\prime}$, at least one of the obedience constraints relative to the deviations highlighted above is binding. Indeed, this would imply that the term on the right-hand side of either inequality 3 or 4 is zero.
We now turn to the conditions under which a particular redistributive transfer is unfeasible for some segmentation. Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$, and assume there exists $p^{\prime},p^{\prime\prime}\in\operatorname{supp}(\sigma_{P})$, $\theta^{\prime}\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p^{\prime\prime})\bigr{)}$ and $\theta^{\prime\prime}\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p^{\prime})\bigr{)}$ with $\theta^{\prime\prime}>\theta^{\prime}\geq p^{\prime\prime}>p^{\prime}$. Consider a redistributive buyer transfer such that a mass $\varepsilon$ of buyers of type $\theta^{\prime}$ are transferred from segment $p^{\prime\prime}$ to $p^{\prime}$, while the same mass $\varepsilon$ of buyers of type $\theta^{\prime\prime}$ are transferred from segment $p^{\prime}$ to $p^{\prime\prime}$, as illustrated in Figure 3(b). This transfer does not impact obedience constraints in segment $p^{\prime}$. Indeed, it decreases the demand at any price $\theta^{\prime}<q\leq\theta^{\prime\prime}$ and does not impact the demand at any other price. In segment $p^{\prime\prime}$, however, the transfer increases the demand at any price $q$ such that $\theta^{\prime}<q\leq\theta^{\prime\prime}$, while leaving the demand at any other price unchanged. Therefore, the redistributive transfer prompts the seller to deviate in segment $p^{\prime\prime}$ to charging a price $\theta^{\prime}<q\leq\theta^{\prime\prime}$ if and only if
$\varepsilon>\frac{1}{q}\Biggl{(}p^{\prime\prime}\sum_{\theta\geq p^{\prime\prime}}\sigma(\theta,p^{\prime\prime})-q\sum_{\theta\geq q}\sigma(\theta,p^{\prime\prime})\Biggr{)}.$ (5)
Hence, no redistributive transfer between types $\theta^{\prime}$ and $\theta^{\prime\prime}$ in segments $p^{\prime}$ and $p^{\prime\prime}$ is feasible for $\sigma$ if, in the upper segment $p^{\prime\prime}$, at least one of the obedience constraints relative to the deviations highlighted above is binding, as this would imply that the term on the right-hand side of inequality 5 is zero.
Proposition 2 establishes necessary and sufficient conditions a segmentation must satisfy for no downward or redistributive transfers to be feasible for that segmentation. To see why, note that condition (ii) guarantees that no buyer of type $\theta$ belonging to some segment $p$ can be transferred to some segment $p^{\prime}$ such that $p<p^{\prime}<\theta$ because the obedience constraint relative to the price $\theta$ is binding in segment $p^{\prime}$. Since redistributive transfers always involve an upward transfer of this sort, condition (ii) ensures that one cannot perform any such transfer. [^21] An upward transfer is defined as the opposite of a downward transfer (Definition 2). Condition (ii) also ensures that some downward transfers are not feasible. Buyers of type $\theta$ cannot be transferred from a segment $p^{\prime}$ to any lower segment $p$ which includes some buyers of type $\theta^{\prime}>\theta$ because the obedience constraint for the price $\theta^{\prime}$ must be binding in segment $p^{\prime}$. However, a downward transfer could in principle be possible if the lower segment $p$ did not include any buyer of type $\theta^{\prime}>\theta$. Condition (i) blocks all these remaining downward transfers. It ensures that even if no buyer of type $\theta^{\prime}>\theta$ is assigned to segment $p$, the obedience constraint relative to some price $\tilde{\theta}\leq\theta$ must be binding in segment $p$, which renders such downward transfer unfeasible.
3.4 Monotone segmentations and progressive pricing
In this subsection, we show that redistributive segmentations exhibit a positive-assortative structure. In other words, redistributive segmentations assign low types to low-price segments and high types to high-price segments. We introduce two notions of positive-assortativity for segmentations.
Definition 4 (Weakly monotone segmentations).
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is weakly monotone if for any $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$,
$\max\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}\leq\max\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p^{\prime})\bigr{)}.$
Definition 5 (Strongly monotone segmentations).
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is strongly monotone if for any $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$,
$\max\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}\leq p^{\prime}.$
Figures 4(a) and 4(b) respectively illustrate weakly and strongly monotone segmentations.
$\theta$$p$
(a) Weak monotonicity.
$\theta$$p$
(b) Strong monotonicity.
Figure 4: Monotone segmentations.
A segmentation is weakly monotone if, for any pair of segments, the highest type assigned to the lower-price segment is not strictly greater than the highest type assigned to the higher-price segment. Whereas a segmentation is strongly monotone if, for any pair of segments, all buyers in the lower-price segment have (weakly) lower types than all buyers in the higher-price segment. [^22] Note that our definition of weak monotonicity implies that the segment supports are weak-set-ordered (Che et al., 2021). Strong monotonicity, on the other hand, implies an ordering of segment supports that is stronger than the strong set order (Topkis, 1998). Note also that, since the intersection of any pair of segment supports contains at most one element, strongly monotone segmentations correspond to a slightly weaker version of monotone partitional information structures (Ivanov, 2021; Mensch, 2021; Onuchic and Ray, 2023; Kolotilin et al., 2024) which partition the state space into convex sets.
Both concepts of monotonicity relate to some form of progressive pricing where high types are charged higher prices than low types. In weakly monotone segmentations, for every buyer of a given type, there exists a buyer of a higher type that is charged a (weakly) higher price. Nevertheless, this notion of progressivity is mild since some buyers of lower type might pay higher prices than some buyers of higher type. This outcome arises in regions where the support of different segments overlap, as buyer types belonging to such overlap are randomly assigned to different segments. On the other hand, strongly monotone segmentations are purely progressive. If a type is consuming at a given price, then every higher type must be assigned to a (weakly) higher segment. Indeed, strongly monotone segmentations assign buyers to segments deterministically, except perhaps for types located at the intersection of two consecutive segments.
A direct implication of condition (ii) in Proposition 2 is that redistributive segmentations are weakly monotone. Indeed, for any $\succsim_{\mathcal{R}}$-maximal segmentation, if a type $\theta$ is in a segment $p\in\operatorname{supp}(\sigma_{P})$, then it must belong to any other segment $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}\leq\theta$, which implies weak monotonicity. [^23] We state the result for strictly redistributive welfare functions to clarify the exposition. If $w\in\mathcal{R}$, then for any $\mu\in\Delta(\Theta)$, there exists a $\mathrm{P}^{\star}_{w,\mu}$-optimal and weakly monotone segmentation of $\mu$.
Theorem 2.A.
For any $\mu\in\Delta(\Theta)$ and $w\in\overline{\mathcal{R}}$, every $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is weakly monotone.
Strongly monotone segmentations, which represent pure progressive pricing, imply that all possible redistributive transfers have been exhausted. This raises the question: under what conditions on redistributive welfare functions do we wish to exhaust all opportunities for implementing such transfers? Suppose we aim to perform a redistributive transfer on a segmentation that is not strongly monotone. While these transfers are always desirable from a welfare perspective, they may not be feasible because they could violate obedience constraints in the upper segment (cf. Figure 3(b)). Nevertheless, it is possible to preserve obedience in the upper segment by transferring a certain mass of the highest type in this segment to a higher-price segment. This series of transfers improves aggregate welfare only if the net welfare gain from the redistributive transfer exceeds the welfare loss due to this upward transfer. The following concept of strongly redistributive welfare functions precisely captures the necessary and sufficient condition for this series of transfers to always be beneficial.
Definition 6 (Strongly redistributive welfare functions).
A welfare function $w\in\overline{\mathcal{R}}$ is strongly redistributive if for any $p,\theta_{k},\theta_{\ell}\in\Theta$ such that $p<\theta_{k}<\theta_{\ell}$,
[…]
A welfare function is strongly redistributive if the net welfare gain from a redistributive transfer involving the two lowest types that could be included in the segment $\theta_{k}$—given by the expression on the left-hand side of condition 6—exceeds by a factor $\theta_{k+1}/(\theta_{k+1}-\theta_{k})$ the maximal welfare loss induced by the upward transfer of any higher type $\theta_{\ell}$—given by the second term on the right-hand side. Figure 5 illustrates the two transfers involved in 6.
$\theta$$p$$p$$\theta_{k}$$\theta_{k}$$\theta_{k+1}$$\theta_{\ell}$$\varepsilon$$\varepsilon$$\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\,\varepsilon$
Figure 5: The combination of a redistributive transfer (in orange) and an upward transfer of high-types in the upper segment (in blue).
As the next result shows, strongly redistributive welfare functions are the only functions for which every optimal segmentation is strongly monotone. We explain condition 6 together with the theorem below.
Theorem 2.B.
A welfare function $w$ is strongly redistributive if and only if for any $\mu\in\Delta(\Theta)$, every $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is strongly monotone.
The proof of Theorem 2.B can be found in Appendix D. In condition 6, $\theta_{k}$ is both the type that benefits from the redistributive transfer and the price charged at the upper segment involved in this transfer. Focusing on such transfers is sufficient because whenever we can implement a redistributive transfer, we can always ensure that it benefits the type equal to the price in the upper segment. Conversely, to exhaust all possible redistributive transfers, it is necessary to perform those that benefit the type corresponding to the price of the upper segment involved in the transfer.
The second term on the right-hand side of 6 corresponds to the maximum welfare loss the type $\theta_{\ell}$ can incur by being transferred upwards from segment $\theta_{k}$. Indeed, the segment $\theta_{\ell}$ is the highest one where the type $\theta_{\ell}$ can be transferred while ensuring efficiency. This welfare loss is scaled by the ratio $\theta_{k+1}/(\theta_{k+1}-\theta_{k})>1$. The reason is the following. If we implement a redistributive transfer of a mass $\varepsilon$ of types $\theta_{k}$ and $\theta_{j}>\theta_{k}$, and this transfer breaks Ob, then the seller will deviate to charging a price $\theta_{k}<p\leq\theta_{j}$ in the upper segment (cf. Figure 3(b)). To preserve obedience, an even larger mass of higher types $\theta_{\ell}\geq\theta_{j}$ must be transferred upward from the upper segment. Specifically, a mass $\varepsilon\times p/(p-\theta_{k})$ of types $\theta_{\ell}$ must be transferred. The ratio $p/(p-\theta_{k})$ decreases with $p$ and is thus maximal for $p=\theta_{k+1}$.
The term on the left-hand side of 6 corresponds to the net welfare gain from a redistributive transfer between two consecutive types. By supermodularity of $w$, the net welfare gain from a redistributive transfer between $\theta_{k}$ and $\theta_{k+1}$ is smaller than the net gain from a redistributive transfer between $\theta_{k}$ and any $\theta_{j}$, with $j>k+1$. As such, a redistributive transfer between these two consecutive types yields the minimal net welfare gain from a redistributive transfer benefiting type $\theta_{k}$.
To summarize, satisfying the inequality 6 implies that the minimal net welfare gains from performing a redistributive transfer benefiting type $\theta_{k}$ outweighs the maximal welfare loss from performing an upward transfer that involves the maximal mass of types $\theta_{\ell}$. Therefore, it is both necessary and sufficient to make any opportunity to implement a redistributive transfer desirable.
We illustrate strongly redistributive welfare functions with Examples 1 and 2. This shows incidentally that the set of these functions is non-empty.
Example 0 (Decreasing Pareto weights, cont’d).
Let $\lambda:\Theta\to\mathbb{R}$ be a strictly decreasing function such that $\lambda(\theta)>0$ for all $\theta\in\Theta$. Define $w$ such that, for any $(\theta,p)\in\Omega$, $w(\theta,p)=\lambda(\theta)(\theta-p)$. By rearranging terms in the inequality 6, we can see that $w$ is strongly redistributive if and only if for any $\theta_{k},\theta_{\ell}\in\Theta$ such that $\theta_{k}<\theta_{\ell}$,
$\frac{\lambda(\theta_{k+1})}{\lambda(\theta_{\ell})}\left(\frac{\lambda(\theta_{k})}{\lambda(\theta_{k+1})}-1\right)\geq\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\frac{\theta_{\ell}-\theta_{k}}{\theta_{k}-\theta_{k-1}}.$ (7)
Inequality 7 imposes a condition on the rate of decrease of $\theta\mapsto\lambda(\theta)$. In this case, we can reinterpret Theorem 2.B as follows: every optimal segmentation is strongly monotone if and only if Pareto weights decrease at a sufficiently fast rate.
Example 1 (Increasing concave transformations, cont’d).
Let $u:\mathbb{R}\to\mathbb{R}$ be strictly increasing and strictly concave. Define $w$ such that, for any $(\theta,p)\in\Omega$, $w(\theta,p)=u(\theta-p)$. In this case, one can show that condition 6 can never be satisfied. Conversely to Example 1, when the redistributive concerns stem from a concave transformation of buyers’ surplus, the welfare gain associated with redistributive transfers can never outweigh the welfare loss associated with burning the surplus of some high type. In this case, Theorem 2.B implies that there always exists a market $\mu\in\Delta(\Theta)$ such that no optimal segmentation is strongly monotone.
Theorem 2.B establishes that a welfare function is strongly redistributive if and only if any optimal segmentation is strongly monotone. Furthermore, as we show in the following lemma (which proof can be found in Appendix E), for any given aggregate market, there is only one strongly monotone segmentation of that market that is $\succsim_{\mathcal{R}}$-maximal. This particular segmentation therefore represents the uniquely optimal market segmentation for any strongly redistributive welfare function.
Lemma 2.
For any $\mu\in\Delta(\Theta)$, there exists a unique strongly monotone segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ that is $\succsim_{\mathcal{R}}$-maximal.
The greedy redistributive algorithm
The simple structure of strongly monotone segmentations and their uniqueness pave the way for a constructive approach. We describe a procedure, called the greedy redistributive algorithm, that outputs, for any market $\mu$, the unique strongly monotone segmentation of $\mu$ that is $\succsim_{\mathcal{R}}$-maximal. We explain in words the functioning of the algorithm and Appendix F contains its pseudo-code.
Strongly monotone segmentations can be seen as the result of lexicographic optimization: it benefits each type sequentially as much as possible, from the smallest to the largest. The greedy redistributive algorithm follows this logic and sequentially assigns types into segments. It starts by creating a lower segment with all buyers of type $\theta_{1}$ and then proceeds to include types $\theta_{2}$ in that segment. If all buyers $\theta_{2}$ can be included while keeping the price in the segment at $\theta_{1}$, the algorithm does so and proceeds to include types $\theta_{3}$. At some point in this process, the algorithm reaches a point such that including all buyers of a given type $\theta_{k}$ would violate Ob in that segment. When the algorithm reaches this point, it fits as many types $\theta_{k}$ as possible in the segment, so that Ob becomes binding at type $\theta_{k}$, and moves on to creating a new segment with the remaining mass of types $\theta_{k}$. This inclusion process is then repeated in the newly created segment starting at type $\theta_{k+1}$, and then proceeds sequentially as previously explained. The algorithm ends when every buyer has been assigned to some segment.
Link to optimal transport
Observe that the optimization problem $\mathrm{P}^{\star}_{w,\mu}$ is similar to the discrete optimal assignment problem (see Galichon, 2018, Chapter 3). The key difference is that instead of optimizing on couplings—the bivariate joint distributions with two prescribed marginals—we optimize on segmentations, which consist of joint distributions such that only the first marginal (the aggregate market $\mu$) is prescribed, while the second marginal (the price distribution $\sigma_{P}$) is endogenous and determined by the seller’s incentive constraints.
It is well established in statistics and optimal transport theory that strongly monotone joint distributions—also known as comonotone distributions—maximize the expectation of supermodular functions over the set of couplings (see, e.g., Theorem 2.1 in Puccetti and Wang, 2015, and Theorem 4.3 in Galichon, 2018). Therefore, since redistributive welfare functions are supermodular, it is natural that redistributive segmentations induce a positive-assortative matching of willingness to pay and prices. Moreover, just as weakly monotone segmentations correspond to the maximal elements of the redistributive order, comonotone distributions correspond to the maximal elements of the supermodular stochastic order (see claim (c) of Theorem 2.1 in Puccetti and Wang, 2015). [^24] As a side note, redistributive transfers (Definition 3) correspond to the mass transfers used in optimal transport to show that comonotone distributions are the extreme points of the transportation polytope (see Peyré and Cuturi, 2019, Proposition 3.4)—i.e. the set of discrete couplings.
Taken together, Propositions 1 and 2 and Theorems 2.A and 2.B thus provide a natural analogue in terms of segmentations to the characterizations of optimal assignments in optimal transport. However, since segmentations are constrained by the seller’s incentives, supermodularity alone no longer ensures comonotonicity. As discussed earlier, our concept of “strong-supermodularity” (Definition 6) becomes a necessary and sufficient condition for achieving comonotone optimal solutions.
3.5 Redistributive rents
We now study how redistributive segmentations distribute surplus between the seller and the buyers. We show that if the welfare function is strongly redistributive, then optimal segmentation might allocate some of the created surplus to the seller. Hence, although the welfare function does not take into consideration the seller’s profit, redistributive segmentations may not maximize total consumer surplus.
For any $\mu\in\Delta(\Theta)$, we say that a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ gives a rent to the seller if the profit under $\sigma$ is strictly higher than the profit under uniform pricing: that is, if
$\sum_{p\in P}p\sum_{\theta\geq p}\sigma(\theta,p)>p^{\star}_{\mu}\sum_{\theta\geq p^{\star}_{\mu}}\mu(\theta).$
The following joint probability distribution over types and prices will prove to be of particular interest. For any market $\mu\in\Delta(\Theta)$, we define $\sigma^{\star}_{\mu}\in\Delta(\Theta\times P)$ as follows:
$\sigma^{\star}_{\mu}(\theta,p)=\left\{\begin{array}[]{ll}\mu(\theta)&\text{if $\theta<p^{\star}_{\mu}$ and $p=\theta_{1}$}\\[15.0pt] \min\left\{\mu(p^{\star}_{\mu}),\displaystyle\frac{\theta_{1}}{p^{\star}_{\mu}-\theta_{1}}\displaystyle\sum_{\theta<p^{\star}_{\mu}}\mu(\theta)\right\}&\text{if $(\theta,p)=(p^{\star}_{\mu},\theta_{1})$}\\[15.0pt] \mu(p^{\star}_{\mu})-\sigma^{\star}_{\mu}(p^{\star}_{\mu},\theta_{1})&\text{if$(\theta,p)=(p^{\star}_{\mu},p^{\star}_{\mu})$}\\[15.0pt] \mu(\theta)&\text{if $\theta>p^{\star}_{\mu}\text{ and }p=p^{\star}_{\mu}$}\\[15.0pt] 0&\text{otherwise.}\end{array}\right.$
[…]
Figure 6: The segmentation $\sigma^{\star}_{\mu}$ for a uniform price $p^{\star}_{\mu}$.
By construction, $\sigma^{\star}_{\mu}$ satisfies $\mathrm{C}_{\mu}$ and $\operatorname{supp}(\sigma^{\star}_{\mu})\subset\Omega$. Therefore, if $\sigma^{\star}_{\mu}$ satisfies Ob then it is a segmentation belonging to $\Sigma_{\Omega}(\mu)$. In this case, we refer to $\sigma^{\star}_{\mu}$ as being feasible. Note that, if $\sigma^{\star}_{\mu}$ is feasible, then it is the unique $\succsim_{\mathcal{R}}$-maximal and strongly monotone segmentation of [^25] Note that $\sigma^{\star}_{\mu}$ is constructed such that, if it is feasible, the seller is indifferent between charging $\theta_{1}$ and $p^{\star}_{\mu}$ in segment $\theta_{1}$. As such, $\sigma^{\star}_{\mu}$ satisfies conditions (i) and (ii) of Proposition 2 if it is feasible, implying that it is $\succsim_{\mathcal{R}}$-maximal. $\mu$. Furthermore, $\sigma^{\star}_{\mu}$ has a very simple structure. It assigns all types lower than $p^{\star}_{\mu}$ to a “discount” segment where the lowest possible price $\theta_{1}$ is charged, and all the remaining types to a residual segment where the uniform price $p^{\star}_{\mu}$ is charged. Figure 6 illustrates $\sigma^{\star}_{\mu}$.
Theorem 3 (Redistributive Rents).
For any $\mu\in\Delta(\Theta)$ and any strongly redistributive welfare function $w$, the two following statements are true.
  1. (i)
    If $\sigma^{\star}_{\mu}$ is feasible, then it is the unique $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation and it does not give any rent to the seller.
  2. (ii)
    If $\sigma^{\star}_{\mu}$ is not feasible, then the unique $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation gives a rent to the seller.
Proof.
A segmentation does not give any rent to the seller if and only if $p^{\star}_{\mu}$ is an optimal price in every market segment. [^26] This follows easily from the fact that the seller’s total profit is the sum of its profits on each segment. Therefore, a strongly monotone segmentation does not give a rent to the seller only if it satisfies two conditions. First, it must generate only two market segments. One containing every type weakly lower than $p^{\star}_{\mu}$, and one containing every type weakly greater than $p^{\star}_{\mu}$. Second, Ob must bind at $p^{\star}_{\mu}$ in the segment $\theta_{1}$. It can easily be seen that $\sigma^{\star}_{\mu}$ is the only strongly monotone segmentation that satisfies those two requirements. Furthermore, following Theorem 2.B, for any strongly redistributive $w$, the $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is the unique strongly monotone one that is $\succsim_{\mathcal{R}}$-maximal. Therefore, if $\sigma^{\star}_{\mu}$ is feasible, then it is uniquely $\mathrm{P}^{\star}_{w,\mu}$-optimal and does not give a rent to the seller. On the contrary, if $\sigma^{\star}_{\mu}$ is not feasible, then the $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation must give a rent to the seller. ∎
Theorem 3 highlights the key tradeoff generated by redistributive concerns. Benefiting poor consumers may require imposing a sacrifice on richer consumers’ surplus. Strongly redistributive welfare functions precisely characterize when the welfare gains from redistribution more than compensate for the surplus loss of higher types. Given that redistributive segmentations are efficient (Lemma 1), as a byproduct of this loss, the seller benefits from a redistributive rent. This result is easily explainable in light of the structure of strongly redistributive segmentations (Theorem 2.B). Such segmentations pool as many low-types as possible in low-price segments, allowing them to benefit from lower prices. Consequently, high types are separated from low types, enabling the seller to better price discriminate. This increases the seller’s profit, except in the special case where $\sigma^{\star}_{\mu}$ is feasible.
4 Implementation through price regulation
Thus far, we have analyzed the segmentation design problem from the perspective of a central planner with redistributive preferences. In practice, however, market segmentation is primarily conducted by private actors with access to consumers’ data, such as platforms. This leads to the problem of implementation: how can a regulator (she) induce the seller (he) to implement a redistributive segmentation?
If the regulator can verify the segmentation chosen by the seller, such an implementation problem is trivial. The regulator can simply commit to punishing the seller if his chosen segmentation differs from her preferred one. In practice, however, it is unlikely that the regulator can perfectly observe the segmentation set by the seller. We show that observing the price distribution induced by a segmentation is sufficient for the regulator to implement her preferred segmentation.
Assume that the regulator and the seller both know the aggregate market $\mu\in\Delta(\Theta)$. The regulator communicates to the seller the segmentation $\sigma$ she wishes to implement, after which the seller chooses a market segmentation $\sigma^{\prime}\in\Sigma(\mu)$, which may or may not coincide with the regulator’s one. The regulator is unable to observe $\sigma^{\prime}$ directly but can verify and enforce punishments based on the induced price distribution [^27] As discussed in the introduction, observing the distribution of prices seems feasible; see Footnote 6. $\sigma^{\prime}_{P}$. Consequently, for any $\mu\in\Delta(\Theta)$, the set of undetectable deviations from $\sigma$ is
$\mathcal{D}_{\mu}(\sigma)\coloneqq\bigl{\{}\sigma^{\prime}\in\Sigma(\mu)\ |\ \sigma^{\prime}_{P}=\sigma_{P}\bigr{\}}.$
Since the regulator can detect any deviation that does not belong to $\mathcal{D}_{\mu}(\sigma)$, she can impose appropriate penalties and deter the seller from deviating to $\sigma^{\prime}\notin\mathcal{D}_{\mu}(\sigma)$. Hence we adopt the following notion of implementability.
Definition 7 (Price-based implementability).
Let $\mu\in\Delta(\Theta)$. A segmentation $\sigma\in\Sigma(\mu)$ is implementable through price-based regulation if there exists no $\sigma^{\prime}\in\mathcal{D}_{\mu}(\sigma)$ such that
$\sum_{p\in\operatorname{supp}(\sigma^{\prime}_{P})}p\sum_{\theta\geq p}\sigma^{\prime}(\theta,p)>\sum_{p\in\operatorname{supp}(\sigma_{P})}p\sum_{\theta\geq p}\sigma(\theta,p).$
A segmentation can be implemented through price-based regulation if there exists no strictly profitable deviation for the seller that is undetectable by the regulator. The next proposition shows that efficiency is, in fact, sufficient for implementability. [^28] Our notion of implementability is equivalent to the notion of credibility defined by Lin and Liu (2024) in the context of Bayesian persuasion. Specifically, the authors define an information structure as credible if there exists no alternative information structure that would strictly increase the sender’s payoff while unaffecting the marginal distribution of signal realizations. As such, Proposition 3 shows that any efficient segmentation is also seller-credible. Our notion of implementation is also equivalent to the notion of price-regulation defined by Schlom (2024) in the context of second-degree price discrimination.
Proposition 3.
Let $\mu\in\Delta(\Theta)$. If $\sigma\in\Sigma_{\Omega}(\mu)$, then $\sigma$ is implementable through price-based regulation.
Proof.
Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$. Recall that, for any $(\theta,p)\in\Omega$ we must have that $\sigma(\theta,p)=\sigma(\theta\,|\,p)\,\sigma_{P}(p)$. We thus obtain the following decomposition of the platform’s profit under $\sigma$:
$\sum_{p\in\operatorname{supp}(\sigma_{P})}p\sum_{\theta\geq p}\sigma(\theta,p)$ $=\sum_{p\in\operatorname{supp}(\sigma_{P})}p\,\sigma_{P}(p)\overbrace{\sum_{\theta\geq p}\sigma(\theta\,|\,p)}^{\mathclap{\text{$=1$ since $\sigma\in\Sigma_{\Omega}(\mu)$}}}$
$=\sum_{p\in\operatorname{supp}(\sigma_{P})}p\,\sigma_{P}(p).$
By definition, $\sum_{\theta\geq p}\sigma^{\prime}(\theta\,|\,p)\leq\sum_{\theta\in\Theta}\sigma^{\prime}(\theta\,|\,p)=1$ for any $\sigma^{\prime}\in\Sigma(\mu)$. Therefore, for any $\sigma^{\prime}\in\mathcal{D}_{\mu}(\sigma)$, we must have in particular that
$\sum_{p\in\operatorname{supp}(\sigma^{\prime}_{P})}p\,\sigma^{\prime}_{P}(p)\sum_{\theta\geq p}\sigma^{\prime}(\theta\,|\,p)$ $\leq\sum_{p\in\operatorname{supp}(\sigma^{\prime}_{P})}p\,\sigma^{\prime}_{P}(p)$
$=\sum_{p\in\operatorname{supp}(\sigma_{P})}p\,\sigma_{P}(p)\,$
This implies that $\sigma$ is implementable through price-based regulation. ∎
Therefore, any efficient segmentation and, following Lemma 1, any redistributive segmentation, can be implemented through price-based regulation. The proof follows from a simple observation. Namely, the difference in the seller’s profit between two segmentations that induce the same price distribution only depends on the mass of consumers that buy the good at each price. For efficient segmentations, this mass is maximal because every buyer consumes. Therefore, there cannot exist any strictly beneficial deviation.
5 Concluding remarks
Price discrimination has long been used as an instrument for redistribution. Many goods and services are provided at reduced prices for disadvantaged groups. For instance, students from low-income backgrounds often pay reduced tuition fees at universities, and it is common for the unemployed or the elderly to receive discounted prices at museums. The prevalence of such redistributive price discrimination indicates that personalized pricing is deemed socially acceptable as long as it is perceived to benefit those in need. This view is supported by experimental data on people’s attitudes to price discrimination (Wu et al., 2012) and is expressed in policy reports (White House Council of Economic Advisers, 2015; Bourreau and De Streel, 2018). Our redistributive segmentations satisfy this social demand by ensuring that price discrimination satisfies a certain degree of price progressivity. Hence, our results inform the current policy debate over the regulation of online personalized pricing. They give qualitative guidelines regarding market segmentations that prioritize consumers with low willingness to pay.
Redistributive price discrimination is usually implemented by public and non-profit organizations (Le Grand, 1975). A concern is that imposing redistributive price discrimination on the private sector could be inefficient. Indeed, conditioning prices on group characteristics could harm profits and deter firms from entering the market (Woodcock, 2022). This paper studies how a private market with a profit-maximizing monopoly can be optimally segmented to achieve a redistributive goal. Redistributive segmentations have the twofold benefit of achieving economic efficiency (Lemma 1) and never harming the seller—Theorem 3 shows that redistributive segmentations can even improve the seller’s profit.
One might suggest that conditioning market segmentation on additional buyer information (e.g., income or group characteristics), rather than only the willingness to pay, would better target redistribution. Nevertheless, as highlighted in Section 4, consumer data is usually collected and processed by private actors for which consumers’ willingness to pay is the payoff-relevant variable. It might also be impossible or illegal for regulators like competition authorities to directly access income data or protected characteristics. Hence, our optimal segmentations constitute a normative benchmark against which to evaluate welfare outcomes that could be achieved by regulating data intermediaries.
Another concern among policymakers is consumers’ privacy (Acquisti et al., 2016; Goldfarb and Que, 2023). The current approach mainly focuses on studying the incentives for consumers to reveal their preferences voluntarily (see Ichihashi, 2020; Hidir and Vellodi, 2020; Ali et al., 2023). Our redistributive segmentations are monotone Theorem 2.A, and thus separate rich and poor consumers in different market segments, sometimes strictly (Theorem 2.B). This separation might conflict with consumer privacy concerns, as it allows the seller to discern buyers’ types better. Studying market segmentations that prioritize poorer consumers’ welfare while satisfying privacy constraints (Strack and Yang, 2024) is a fruitful avenue for future research.
Appendix A Proof of Theorem 1
Proof.
For any $f,g\in\mathbb{R}^{\Omega}$ we let $\langle f,g\rangle$ be defined as
$\langle f,g\rangle=\sum_{(\theta,p)\in\Omega}f(\theta,p)\,g(\theta,p).$
We know from Lemma 1 that we can, without loss of generality, restrict attention to segmentations supported on $\Omega$. Hence, we treat a segmentation $\sigma$ as an element of $\mathbb{R}^{\Omega}$.
We first define the following set of $\mathcal{E}$ is the set of real functions $w$ defined on $\Omega$ that is supermodular and such that, for any $\theta\in\Theta$, the function $p\mapsto w(\theta,p)$ is non-increasing on $\{\theta_{1},\dots,\theta\}$.
=$\mathcal{E}=\Bigl{\{}v\in\mathbb{R}^{\Omega}\ \big{|}\ \exists w\in\mathcal{R}\text{ and }b\in\mathbb{R}\text{ such that},\forall(\theta,p)\in\Omega,\,v(\theta,p)=w(\theta,p)+b\Bigr{\}}.$
For any given aggregate market $\mu$, we define the order $\succsim_{\mathcal{E}}$ over the set $\Sigma_{\Omega}(\mu)$ in the analogous way we defined the order $\succsim_{\mathcal{R}}$. The following lemma allows us to work with the family of functions $\mathcal{E}$.
Lemma 3.
$\succsim_{\mathcal{E}}=\succsim_{\mathcal{R}}$.
Proof.
As $\mathcal{R}\subset\mathcal{E}$, if $\sigma\succsim_{\mathcal{E}}\sigma^{\prime}$, then $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$. Conversely, let $\sigma,\sigma^{\prime}$ be such that $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$, that is, for any $w\in\mathcal{R}$, $\langle\sigma,w\rangle\geq\langle\sigma^{\prime},w\rangle$. This implies that for any $w\in\mathcal{R}$ and $b\in\mathbb{R}$, $\langle\sigma,w\rangle+b\geq\langle\sigma^{\prime},w\rangle+b$. With a slight abuse of notation, by denoting $b$ also as the element of $\mathbb{R}^{\Theta\times P}$ constantly equal to $b$, $\langle\sigma,b\rangle=\langle\sigma^{\prime},b\rangle=b$. Therefore, for any $w\in\mathcal{R}$ and $b\in\mathbb{R}$,
$\langle\sigma,w\rangle+b\geq\langle\sigma^{\prime},w\rangle+b$ $\iff\langle\sigma,w\rangle+\langle\sigma,b\rangle\geq\langle\sigma^{\prime},w\rangle+\langle\sigma^{\prime},b\rangle$
$\iff\langle\sigma,w+b\rangle\geq\langle\sigma^{\prime},w+b\rangle$
This implies that $\sigma\succsim_{\mathcal{E}}\sigma^{\prime}$. ∎
We now prove the statement of the theorem for the order $\succsim_{\mathcal{E}}$. The proof follows similar steps as the proof of Theorem 1 in Meyer and Strulovici (2015). We first show that the set $\mathcal{E}$ is the dual cone of the set $\mathcal{T}$, as shown by the next lemma.
Lemma 4.
For any $w\in\mathbb{R}^{\Omega}$,
$w\in\mathcal{E}\iff\forall t\in\mathcal{T},\;\langle w,t\rangle\geq 0.$
Proof.
$\Rightarrow$ This follows from the definition of $\mathcal{R}$ and therefore of $\mathcal{E}$.
$\Leftarrow$ Let $w$ be such that $\langle w,t\rangle\geq 0$ for every $t\in\mathcal{T}$. By contradiction, suppose there exists $(\theta,p),(\theta,p^{\prime})\in\Omega$ such that $p<p^{\prime}$ and $w(\theta,p)<w(\theta,p^{\prime})$. Take the downward buyer transfer $d\in\mathbb{R}^{\Omega}$ such that $d(\theta,p)=-d(\theta,p^{\prime})=1$, then $\langle w,d\rangle<0$, a contradiction. Similarly, suppose there exists $(\theta,p),(\theta^{\prime},p^{\prime})\in\Omega$ such that $\theta<\theta^{\prime}$, $p<p^{\prime}$, and $w(\theta,p)-w(\theta,p^{\prime})<w(\theta^{\prime},p)-w(\theta^{\prime},p^{\prime})$. In that case, take the redistributive buyer transfer $r\in\mathbb{R}^{\Omega}$ such that
$r(\theta,p)=r(\theta^{\prime},p^{\prime})=-r(\theta^{\prime},p)=-r(\theta,p^{\prime})=1,$
then $\langle w,r\rangle<0$, again a contradiction. ∎
Recall that the conic hull of $\mathcal{T}$ is the smallest convex cone containing $\mathcal{T}$ and equals the set of all positive linear combinations of finitely many elements of $\mathcal{T}$ (see Ok, 2007, p. 430):
$\mathrm{cone}(\mathcal{T})=\left\{\sum_{t\in T}\lambda_{t}t\ \Big{|}\ \text{$T\subset\mathcal{T}$, $|T\rvert<\infty$ and $\lambda\in\mathbb{R}_{+}^{T}$}\right\}.$
Note that for any $t\in\mathcal{T}$ and $\lambda_{t}\geq 0$, $\lambda_{t}t\in\mathcal{T}$, therefore we can define $\mathrm{cone}(\mathcal{T})$ as
$\mathrm{cone}(\mathcal{T})=\left\{\sum_{i=1}^{n}t_{i}\ \Big{|}\ \text{$n\in\mathbb{N}^{*}$ and $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$}\right\}.$
As $\mathrm{cone}(\mathcal{T})$ is closed, we can show in the next lemma that it is the dual cone of $\mathcal{E}$ (see Luenberger, 1969, p.215).
Lemma 5.
For any $\tau\in\mathbb{R}^{\Omega}$,
$\tau\in\mathrm{cone}(\mathcal{T})\iff\forall w\in\mathcal{E},\;\langle w,\tau\rangle\geq 0.$
Proof.
$\Rightarrow$ Assume $\tau\in\mathrm{cone}(\mathcal{T})$. Hence, there exists $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that $\tau=\sum_{i=1}^{n}t_{i}$. Following Lemma 4 we have that, for any $w\in\mathcal{E}$,
$\langle\tau,w\rangle=\sum_{i=1}^{n}\langle t_{i},w\rangle\geq 0.$
$\Leftarrow$ Assume $\upsilon\notin\mathrm{cone}(\mathcal{T})$. As $\mathrm{cone}(\mathcal{T})$ is closed, the separating hyperplane theorem implies that there exists $w\in\mathbb{R}^{\Omega}$ such that $\langle\upsilon,w\rangle<\langle\tau,w\rangle$ for every $\tau\in\mathrm{cone}(\mathcal{T})$. Since $\mathrm{cone}(\mathcal{T})$ is a cone, it must be that $\langle\tau,w\rangle\geq 0$ for any $\tau\in\mathrm{cone}(\mathcal{T})$ as otherwise there would exist some $\alpha>0$ such that $\langle\alpha\tau,w\rangle\leq\langle\upsilon,w\rangle$. Thus, Lemma 4 implies that $w\in\mathcal{E}$. Finally, since $0=\min_{\tau\in\mathrm{cone}(\mathcal{T})}\langle\tau,w\rangle$, we must have $\langle\upsilon,w\rangle<0$. ∎
We can now restate the theorem as follows: for any $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$, $\sigma\succsim_{\mathcal{E}}\sigma^{\prime}$ if and only if there exists $n\in\mathbb{N}^{*}$ and a sequence $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that:
$\sigma=\sigma^{\prime}+\sum_{i=1}^{n}t_{i}$ (8)
Equation 8 is equivalent to $\sigma-\sigma^{\prime}\in\mathrm{cone}(\mathcal{T})$, and by Lemma 5, this is equivalent to $\langle\sigma-\sigma^{\prime},w\rangle\geq 0$ for every $w\in\mathcal{E}$. ∎
Appendix B Proof of Proposition 1
Claim (i) was proven in the main text. We now prove claim (ii).
Proof.
For any set $A$ we let $-A=\bigl{\{}-a\ |\ a\in A\bigr{\}}$. For any two sets $A$ and $B$, we let $A+B\coloneqq\bigl{\{}a+b\ |\ a\in A,b\in B\bigr{\}}$ be the Minkowski sum of $A$ and $B$, and we let $A-B=A+(-B)$.
The set $\Sigma_{\Omega}(\mu)$ is a convex polytope. Let $\sigma\in\Sigma_{\Omega}(\mu)$ be $\succsim_{\mathcal{R}}$-maximal. Theorem 1 implies that for any $t\in\mathrm{cone}(\mathcal{T})$ such that $t\neq 0$, we must have $\sigma+t\notin\Sigma_{\Omega}(\mu)$. Therefore, $\bigl{[}\Sigma_{\Omega}(\mu)\cap\bigl{(}\{\sigma\}+\mathrm{cone}(\mathcal{T})\bigr{)}\bigr{]}=\{\sigma\}$. This implies that $\sigma$ is on the relative border of both $\Sigma_{\Omega}(\mu)$ and $\{\sigma\}+\mathrm{cone}(\mathcal{T})$. Given that these are closed convex sets, the separating hyperplane theorem implies that there exists $v\in\mathbb{R}^{\Omega}$ such that:
$\forall\sigma^{\prime}\in\Sigma_{\Omega}(\mu),\forall t\in\mathrm{cone}(\mathcal{T}),\quad\langle\sigma^{\prime},v\rangle\leq\langle\sigma,v\rangle\leq\langle\sigma+t,v\rangle.$
Thus, for any $t\in\mathrm{cone}(\mathcal{T})$,
$\langle\sigma,v\rangle\leq\langle\sigma+t,v\rangle\iff 0\leq\langle t,v\rangle,$
which, by Lemma 4, implies that $v\in\mathcal{E}$. Therefore, there exists $w\in\mathcal{R}$ and $b\in\mathbb{R}$ such that $v=w+b$. Hence, for any $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$,
$\langle\sigma^{\prime},v\rangle\leq\langle\sigma,v\rangle\iff\langle\sigma^{\prime},w+b\rangle\leq\langle\sigma,w+b\rangle\iff\langle\sigma^{\prime},w\rangle\leq\langle\sigma,w\rangle.$
Therefore, $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal. ∎
Appendix C Proof of Proposition 2
From Theorem 1, for any $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$, there exists no $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ such that $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$ if and only if for any $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tau\neq 0$, $\sigma+\tau\notin\Sigma_{\Omega}(\mu)$. We first show the following equivalence.
Lemma 6.
Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$. For any $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tau\neq 0$, $\sigma+\tau\notin\Sigma_{\Omega}(\mu)$ if and only if for any $t\in\mathcal{T}$ such that $t\neq 0$, $\sigma+t\notin\Sigma_{\Omega}(\mu)$.
Proof.
Necessity is straightforward. Let now $\sigma\in\Sigma_{\Omega}(\mu)$ be such that for any $t\in\mathcal{T}$ such that $t\neq 0$, $\sigma+t\notin\Sigma_{\Omega}(\mu)$. First, this means that $\sigma$ is on the relative border of $\Sigma_{\Omega}(\mu)$, as otherwise, for any neighbourhood around $\sigma$ included in $\Sigma_{\Omega}(\mu)$, there would exist sufficiently small transfers $t\in\mathcal{T}$ such that $\sigma+t$ is in this neighbourhood. Second, this also implies that $\sigma\neq\sigma^{F}$, where $\sigma^{F}$ is the first-degree price discrimination segmentation. That is, the segmentation that allocates each buyer to the segment where the price equals their type. This is because downward transfers could be performed on $\sigma^{F}$ without violating the obedience constraints.
We now proceed by contradiction, by assuming that there exists $\tau\in\mathrm{cone}(\mathcal{T})$, $\tau\neq 0$, such that $\sigma^{\prime}\coloneqq\sigma+\tau\in\Sigma_{\Omega}(\mu)$, that is, $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$. As in Appendix B, we let $-A=\bigl{\{}-a\ |\ a\in A\bigr{\}}$ for any set $A$, and $A+B\coloneqq\bigl{\{}a+b\ |\ a\in A,b\in B\bigr{\}}$ and $A-B=A+(-B)$ for any two sets $A$ and $B$. We first prove that $\Sigma_{\Omega}(\mu)\subset\bigl{(}\{\sigma\}+\mathrm{cone}(\mathcal{T})\bigr{)}$, which is equivalent to $\bigl{(}\Sigma_{\Omega}(\mu)-\{\sigma\}\bigr{)}\subset\mathrm{cone}(\mathcal{T})$. Given that $\sigma^{\prime}=\sigma+\tau\in\Sigma_{\Omega}(\mu)$, for some $\tau\neq 0$ and $\tau\notin\mathcal{T}$ this means that $\bigl{(}\Sigma_{\Omega}(\mu)-\{\sigma\}\bigr{)}\,\cap\,\operatorname{int}\bigl{(}\mathrm{cone}(\mathcal{T})\bigr{)}\neq\varnothing$. Now suppose by contradiction that $\bigl{(}\Sigma_{\Omega}(\mu)-\{\sigma\}\bigr{)}\not\subset\mathrm{cone}(\mathcal{T})$. As $\Sigma_{\Omega}(\mu)$ and $\mathrm{cone}(\mathcal{T})$ are closed and convex sets, there exists $x\in\Sigma_{\Omega}(\mu)$ that is on relative border of $\mathrm{cone}(\mathcal{T})$. This means that there exists $\tilde{\sigma}\in\Sigma_{\Omega}(\mu)$ such that $\tilde{\sigma}=\sigma+t$ for some $t\in\mathcal{T}$ (as the borders of $\mathrm{cone}(\mathcal{T})$ are elements of $\mathcal{T}$), which contradicts our premise. Hence we proved that $\Sigma_{\Omega}(\mu)-\{\sigma\}\subset\mathrm{cone}(\mathcal{T})$; which implies that for any $\tilde{\sigma}\in\Sigma_{\Omega}(\mu)$, there exists $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tilde{\sigma}=\sigma+\tau$. To see that this is a contradiction, consider $\sigma^{F}$. We know that $\sigma^{F}\in\Sigma_{\Omega}(\mu)$ and $\sigma^{F}\neq\sigma$, but $\sigma^{F}$ cannot be reached from $\sigma$ by a finite sequence of transfers in $\mathcal{T}$; that is, there exists no $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\sigma^{F}=\sigma+\tau$. ∎
Equipped with Lemma 6, we can now proceed to the proof of Proposition 2 by proving it is equivalent that a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ satisfies claims (i) and (ii) in the proposition, and that there exists no $t\in\mathcal{T}$ such that $t\neq 0$ and $\sigma+t\in\Sigma_{\Omega}(\mu)$. For any $\mu\in\Delta(\Theta)$, $\sigma\in\Sigma_{\Omega}(\mu)$ and $p\in\operatorname{supp}(\sigma_{P})$, define:
$\mathcal{I}_{\sigma}(p)=\Biggl{\{}q\in\Theta\setminus\{p\}\;\Big{|}\;p\sum_{\theta\geq p}\sigma(\theta,p)=q\sum_{\theta\geq\tilde{p}}\sigma(\theta,q)\Biggr{\}}.$
Proof of Proposition 2.
$\Rightarrow$ Suppose that that there exists no $t\in\mathcal{T}$ such that $t\neq 0$ and $\sigma+t\in\Sigma_{\Omega}(\mu)$. Note first that downward and redistributive buyer transfers do not change the marginal distribution over types. Hence, for any $\sigma^{\prime}$ obtained from $\sigma$ by such transfers, we can restrict our attention to the case where $\sigma^{\prime}\notin\Sigma_{\Omega}(\mu)$ is caused by $\sigma^{\prime}$ violating Ob. We first prove that claim (i) in Proposition 2 is satisfied. Denote $\bar{p}=\max\operatorname{supp}(\sigma_{P}))$ and $\bar{\theta}=\max\bigl{\{}\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,\bar{p})\bigr{)}\bigr{\}}$, and let $p\in\operatorname{supp}(\sigma_{P})$ be such that $p<\bar{p}$. As there exists no $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ that is obtained from $\sigma$ by a downward buyer transfer, this means that there exists $p<q\leq\bar{\theta}$ such that $q\in\mathcal{I}_{\sigma}(p)$, as otherwise, we could construct $\sigma^{\prime}$ from $\sigma$ by a downward buyer transfer of mass $\varepsilon>0$ from $(\bar{\theta},\bar{p})$ to $(\bar{\theta},p)$ such that $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ for $\varepsilon$ sufficiently small. This implies that $|\mathcal{I}_{\sigma}(p)|\geq 2$.
We now prove claim (ii). Let $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$. By contradiction, assume there exists $q\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}$ such that $q\geq p^{\prime}$ and that $q\notin\mathcal{I}_{\sigma}(p^{\prime})$. This implies that $q>p^{\prime}$, so that the greatest type in $\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}$ that is strictly lower than $q$, defined by
$\theta_{q}:=\max\Bigl{\{}\theta\in\Theta\ \big{|}\ \theta<q,\,\theta\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}\Bigr{\}},$
is well-defined. Then, for $\varepsilon>0$ sufficiently small, the segmentation that is obtained from $\sigma$ by the redistributive buyer transfer that transfers mass $\varepsilon$ from $(\theta_{q},p^{\prime})$ and $(q,p)$ to $(\theta_{q},p)$ and $(q,p^{\prime})$ satisfies Ob, a contradiction.
$\Leftarrow$ Suppose that $\sigma\in\Sigma_{\Omega}(\mu)$ satisfies points (i) and (ii) in Proposition 2. We first show that there exists no $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ that can be obtained from $\sigma$ by a redistributive buyer transfer. By contradiction, this would imply a segmentation $\sigma^{\prime}$ obtained from $\sigma$ by a transfer of an $\varepsilon$ mass from $(\theta^{\prime},p^{\prime\prime})$ and $(\theta^{\prime\prime},p^{\prime})$ with $p^{\prime}<p^{\prime\prime}$ and $\theta^{\prime}<\theta^{\prime\prime}$ to $(\theta^{\prime},p^{\prime})$ and $(\theta^{\prime\prime},p^{\prime\prime})\in\operatorname{supp}(\sigma)$. Claim (ii) in Proposition 2 implies that $\theta^{\prime\prime}\in\mathcal{I}_{\sigma}(p^{\prime\prime})$, which in turn means that such a $\sigma^{\prime}$ cannot satisfy Ob.
We now show that there exists no $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$ that can be obtained from $\sigma$ by a downward buyer transfer. By contradiction take $(\theta^{\prime},p^{\prime\prime})\in\operatorname{supp}(\sigma)$, $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p^{\prime}<p^{\prime\prime}$, $\varepsilon>0$, and suppose that $\sigma^{\prime}$ obtained from $\sigma$ by a downward buyer transfer from $(\theta^{\prime},p^{\prime\prime})$ to $(\theta^{\prime},p^{\prime})$ is in $\Sigma_{\Omega}(\mu)$. Claim (i) in Proposition 2 implies that there exists $q>p^{\prime}$ such that $q\in\mathcal{I}_{\sigma}(p^{\prime})$. If $q\leq\theta^{\prime}$, then we must have
$p^{\prime}\sum_{\theta\geq p^{\prime}}\sigma^{\prime}(\theta,p^{\prime})<q\sum_{\theta\geq q}\sigma^{\prime}(\theta,p^{\prime}),$
which contradicts $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$. Therefore, it must be that $q>\theta^{\prime}\geq p^{\prime\prime}$. Now, claim (ii) in Proposition 2 implies that $q\in\mathcal{I}_{\sigma}(p^{\prime\prime})$, meaning that
$p^{\prime\prime}\sum_{\theta\geq p^{\prime\prime}}\sigma^{\prime}(\theta,p^{\prime\prime})<p^{\prime\prime}\sum_{\theta\geq p^{\prime\prime}}\sigma(\theta,p^{\prime\prime})=q\sum_{\theta\geq q}\sigma(\theta,p^{\prime\prime})=q\sum_{\theta\geq q}\sigma^{\prime}(\theta,p^{\prime\prime}),$
which contradicts $\sigma^{\prime}\in\Sigma_{\Omega}(\mu)$. ∎
Appendix D Proof of Theorem 2.B
Proof of claim (ii).
$\Leftarrow$ Consider a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ that is $\succsim_{\mathcal{R}}$-maximal but not strongly monotone. Thus, there exists at least two segments $p,\theta_{i}\in\operatorname{supp}(\sigma_{P})$ such that $\max\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}>\theta_{i}$. We denote as $\theta^{\prime}$ the lowest type in the segment $p$ that is strictly above $\theta$, and as $\theta_{k}$ the maximal type in the segment $\theta_{i}$:
$\theta^{\prime}=\min\bigl{\{}x\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p)\bigr{)}\ |\ x>\theta_{i}\bigr{\}},$
$\theta_{k}=\max\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,\theta_{i})\bigr{)}.$
Let $\varepsilon>0$ be sufficiently small and $r$ be the redistributive buyer transfer such that
$r(\theta_{i},p)=r(\theta^{\prime},\theta_{i})=-r(\theta_{i},\theta_{i})=-r(\theta^{\prime},p)=\varepsilon,$
and $r=0$ otherwise. Neglecting for the moment the negative indirect impacts of $r$ by triggering a violation of the obedient constraint in the segment $\theta_{i}$, the direct impact on the welfare of types $\theta_{i}$ and $\theta^{\prime}$ involved in the transfer $r$ is
$\varepsilon\left[(w(\theta_{i},p)-w(\theta_{i},\theta_{i}))-\left(w(\theta^{\prime},p)-w(\theta^{\prime},\theta_{i})\right)\right]$ (9)
$>$ $\varepsilon\left[(w(\theta_{i},p)-w(\theta_{i},\theta_{i}))-\left(w(\theta_{i+1},p)-w(\theta_{i+1},\theta_{i})\right)\right]>0,$ (10)
where both strict inequalities come from $w\in\overline{\mathcal{R}}$. In addition, given that $\sigma$ is $\succsim_{\mathcal{R}}$-maximal, $\sigma+r$ does not satisfy Ob. In particular, there exists $\hat{\theta}$, with $\theta_{i}<\hat{\theta}\leq\theta^{\prime}$, such that $\hat{\theta}\in\mathcal{I}^{\sigma}(\theta_{i})$ (see Appendix C for the definition of $\mathcal{I}^{\sigma}(\theta)$), and thus:
$\theta\sum_{x\geq\theta}\sigma(x,\theta)<\hat{\theta}\left(\sum_{x\geq\hat{\theta}}\sigma(x,\theta)+\varepsilon\right).$
To recover the obedience of the price recommendation $\theta_{i}$, we implement the transfer $t\colon\Omega\to\mathbb{R}$, which transfers a mass $\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon$ of buyers from $(\theta_{k},\theta_{i})$ to $(\theta_{k},\theta_{k})$: that is,
$t(\theta_{k},\theta_{k})=-t(\theta_{k},\theta_{i})=\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon,$
and $t=0$ otherwise. First, for sufficiently small $\varepsilon$, $\sigma+t\in\Delta(\Omega)$. Second, implementing $t$ cannot break any obedience constraints, and it corrects from any violation of obedience constraints induced by the transfer $r$. To see this, look at the difference between the profit made by setting the price $\theta_{i}$ and setting the price $q\geq\theta_{i+1}$:
$\theta_{i}\left(\sum_{x\geq\theta_{i}}\sigma(x,\theta_{i})-\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)-q\left(\sum_{x\geq q}\sigma(x,\theta_{i})+\varepsilon-\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$
$\geq$ $\varepsilon\left(\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}(q-\theta_{i})-q\right)=\varepsilon\frac{\theta_{i}(q-\theta_{i+1})}{\theta_{i+1}-\theta_{i}}\geq 0$
where the first inequality follows from $\theta_{i}$ being an optimal price in segment $\theta_{i}$ according $\sigma$, and the second inequality follows $q\geq\theta_{i+1}$.
Therefore, $\sigma+r+t\in\Sigma_{\Omega}(\mu)$. The loss from the transfer $t$ is
$\frac{\theta_{i+1}}{\theta_{i+1}-\theta}\varepsilon\bigr{[}w(\theta_{k},\theta_{i})-w(\theta_{k},\theta_{k})\bigr{]}.$ (11)
Therefore, implementing the transfers $r$ and $t$ is beneficial if and only if 9 $\geq$ 11. If $w$ is strongly redistributive, we have that 10 $\geq$ 11. Given that 9 $\geq$ 10, this ends the proof.
$\Rightarrow$ We prove the contrapositive. Suppose that $w\in\overline{\mathcal{R}}$ is not strongly redistributive. So there exist $p<\theta_{i}<\theta_{i+1}\leq\theta_{k}$, such that
$\bigl{[}w(\theta_{i},p)-w(\theta_{i},\theta_{i})\bigr{]}-\bigl{[}w(\theta_{i+1},p)-w(\theta_{i+1},\theta_{i})\bigr{]}<\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\bigl{[}w(\theta_{k},\theta_{i})-w(\theta_{k},\theta_{k})\bigr{]}.$ (12)
Let $\mu\in\Delta\bigl{(}\{p,\theta_{i},\theta_{i+1},\theta_{k}\}\bigr{)}$ such that the unique strongly monotone and $\succsim_{\mathcal{R}}$-maximal segmentation, denoted $\sigma$, has three segments:
  • one segment $p$ including all types $p$ and some type $\theta_{i}$;
  • one segment $\theta_{i}$ including the remaining types $\theta_{i}$ and all types $\theta_{i+1}$;
  • and one segment $\theta_{k}$ including all types $\theta_{k}$.
It is easy to show that such a $\mu$ exists. [^30] It can be provided upon requests. Let $\varepsilon>0$ be sufficiently small. We will now perform two transfers. The first one is a downward transfer $d$ defined by:
$d(\theta_{k},\theta_{i})=-d(\theta_{k},\theta_{k})=\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon,$
$d=0$ otherwise. The second transfer, $s$, is the reverse of a redistributive one. It is defined by:
$s(\theta_{i},\theta_{i})=s(\theta_{i+1},p)=-s(\theta_{i+1},\theta_{i})=-s(\theta_{i},p)-\varepsilon,$
and $s=0$ otherwise. We check that $\sigma+d+s$ satisfies Ob. First, for sufficiently small $\varepsilon$, this does not change the optimality of price $p$ in segment $p$. Second, in segment $\theta_{i}$, the transfers do not change the obedience constraint between for $\theta_{i+1}$ as
$\theta_{i}\left(\sigma(\theta_{i},\theta_{i})+\sigma(\theta_{i},\theta_{i+1})+\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$ $-\theta_{i+1}\left(\sigma(\theta_{i},\theta_{i+1})-\varepsilon+\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$
$=$ $\varepsilon\left(\frac{\theta_{i}\theta_{i+1}}{\theta_{i+1}-\theta_{1}}-\frac{\theta_{i}\theta_{i+1}}{\theta_{i+1}-\theta_{1}}\right)$
$=$ $0,$
where the first equality follows from $\theta_{i+1}$ being an optimal price in segment $\theta_{i}$ for segmentation $\sigma$. Similarly, one can show that charging $\theta_{i}$ always gives a greater profit than charging $\theta_{k}$ if the segmentation is $\sigma+d+s$. Therefore $\sigma+d+s\in\Sigma_{\Omega}(\mu)$. Furthermore, it directly follows from Equation 12 that $\langle w,\sigma+d+s\rangle>\langle w,\sigma\rangle$. Therefore, $\sigma$, which is the unique strongly monotone and $\succsim_{\mathcal{R}}$-maximal segmentation of $\mu$ (Lemma 2) is not optimal for $w$. ∎
Appendix E Proof of Lemma 2
Proof.
First, if the uniform price of $\mu$ is $\theta_{1}$, there exists unique segmentation that is $\succsim_{\mathcal{R}}$-maximal, the one that does not segment the market. Suppose now that the uniform price of $\mu$ is strictly greater than $\theta_{1}$. Let $\sigma\in\Sigma_{\Omega}(\mu)$ be strongly monotone and $\succsim_{\mathcal{R}}$-maximal. Therefore it satisfies the conditions of Proposition 2. Denote as $\theta_{k}$ the maximal type that is in the segment $\theta_{1}$. By Proposition 2, there exists $\theta_{i}$, with $\theta_{1}<\theta_{i}\leq\theta_{k}$ such that $\theta_{i}\in\mathcal{I}_{\sigma}(\theta_{1})$. Consequently, any segmentation that is strongly monotone and $\succsim_{\mathcal{R}}$-maximal must have the same market segment $\theta_{1}$. By iterating the same argument for the other segments (starting with the next segment $\theta_{k}$ or $\theta_{k+1}$), this implies that $\sigma$ is the unique segmentation that satisfies these conditions. ∎
Appendix F Pseudo-code for the greedy redistributive algorithm
Algorithm 1 Greedy redistributive algorithm
$\sigma\leftarrow 0_{K\times K}$; $\triangleright$ Initiates the segmentation as being the null matrix of size $K\times K$. $\sigma_{i,j}$ denotes the element of $\sigma$ in row $i$ and column $j$ and $\sigma_{i,\bullet}$ denotes the row $i$ of $\sigma$ seen as a vector in $\mathbb{R}^{K}$.
[…]
$\delta_{k}\leftarrow\delta_{k}+x$;
[…]
$\sigma_{p,t}\leftarrow m$; $\triangleright$ Allocate this mass to segment $p$.
$p\leftarrow t$; $\triangleright$ Initiate a new segment with price $t$.
[…]