### v5

Abstract
We study how to optimally segment monopolistic markets with a redistributive objective. We characterize optimal redistributive segmentations and show that they (i) induce the seller to price progressively, i.e., charge richer consumers higher prices than poorer ones, and (ii) may not maximize consumer surplus, instead granting extra profits to the monopolist. We further show that optimal redistributive segmentations are implementable via price-based regulation.
JEL classification codes: D42, D83, D39.
Keywords: Market Segmentation; Price Discrimination; Redistribution; Information Design.
1. Introduction
Firms and online platforms use rich consumer data to segment markets and personalize prices. [^1] Third-degree price discrimination is often implemented through targeted discounts. For instance, this is how Amazon Brand Tailored promotions and Google promotion assets operate. See mohammed17how; wallheimer18are; oecd2022personalised for more details and evidence. While these practices can improve economic efficiency and even increase total consumer surplus, questions remain about their distributional consequences.
Distributive concerns have long shaped both policy practice and fairness norms regarding price discrimination. From “lifeline” tariffs to student discounts at cultural activities or food stamps to low-income households, price discrimination has repeatedly been justified on the basis of who it benefits. Recent regulatory debates on online personalized pricing have expressed similar concerns (WhiteHouse2015; bourreau2018regulation). Despite this, most of the theoretical literature on price discrimination has focused on studying its impacts on aggregate measures such as output or total consumer surplus. Price discrimination inherently affects consumers differently, and evaluating its welfare effects solely through such aggregates conceals these heterogeneous impacts.
In particular, segmenting markets to maximize total consumer surplus can favor wealthier consumers. As bbm15 show, monopolistic markets can be segmented so that the whole generated surplus accrues to consumers. This outcome is achieved by exploiting pooling externalities (galperti2023): consumers with a low willingness to pay exert a positive externality on consumers with a high willingness to pay who belong to the same segment by inducing the seller to charge prices below the uniform monopoly price. Adequately pooling consumers in the same segments therefore increases the total share of the surplus going to consumers, particularly those most willing to pay. However, a low willingness to pay is often indicative of adverse economic and social circumstances. Therefore, maximizing consumer surplus can benefit wealthy buyers to the detriment of poorer ones.
This paper aims to better understand these distributive effects of market segmentation. Specifically, we address the normative question of how to segment a market to benefit consumers while prioritizing the welfare of those with low willingness to pay. Our two main economic insights are the following. First, unlike segmentations that maximize total consumer surplus, which assign consumers with a high willingness to pay to low-price segments, optimal redistributive segmentations are positive-assortative: they assign buyers with a high willingness to pay to high-price segments and buyers with low willingness to pay to low-price segments. As a result, they induce the monopolist to price progressively, meaning that wealthier consumers pay higher prices than poorer ones (Section˜3.3). Second, this separation of wealthy and poor consumers in different segments sometimes allows the seller to price discriminate more effectively, and therefore increase its profit compared to uniform pricing. Hence, redistributive segmentations may not maximize total consumer surplus (Section˜3.4).
Our model considers a continuous population of buyers, each having a unit demand for a good sold by a monopolistic seller. Every buyer is characterized by her willingness to pay for the good—her type. We refer to the distribution of types in the economy as the aggregate market. Without loss of generality, we assume that the seller’s cost of production is zero, so gains from trade are always positive. A market designer can flexibly segment the aggregate market, anticipating that the seller will charge a profit-maximizing price in each segment.
The designer’s preferences are represented by a welfare function, which describes the social value of a given type consuming at a given price. We focus on welfare functions that align with consumer surplus: welfare equals zero when buyers do not purchase, and is decreasing in price when buyers purchase. Within this class, we study redistributive welfare functions, for which the welfare gain resulting from reducing the price paid by buyers with low type is greater than the welfare gain resulting from an equal price reduction for buyers with high type. [^2] Mathematically, this corresponds to the welfare function being supermodular. This is motivated by the idea that buyers’ willingness to pay reflects both their taste for the good and their opportunity cost of money. Prioritizing low willingness to pay buyers thus appropriately captures redistributive concerns when heterogeneity in willingness to pay primarily reflects differences in ability to pay: wealthier individuals can typically pay more for what they need. [^3] We refer to dworczak21; akbarpour2024redistributive; dworczak2024inequality for in-depth discussions of this argument. Examples of redistributive welfare functions include weighted consumer surplus with Pareto weights decreasing in willingness to pay, or increasing concave transformations of consumer surplus.
Our main result, Theorem˜1, characterizes the set of segmentations that are optimal for some redistributive welfare function as those that saturate a specific set of the seller’s incentive constraints. These binding incentive constraints ensure that no marginal reallocations of buyers across segments can be beneficial for every redistributive welfare function. Theorem˜1 relies on a key result of independent interest. We define a redistributive order, which ranks a segmentation as more redistributive than another if it yields higher aggregate welfare for every redistributive welfare function. We characterize it in terms of elementary reallocations of buyers across segments: a segmentation is more redistributive than another if and only if there exists a sequence of such elementary reallocations transforming the latter into the former (Proposition˜1).
We leverage Theorem˜1 to draw key economic implications of optimal redistributive segmentations. We first show the optimality of progressive pricing by proving that any optimal segmentation must assign buyer types to prices in a positive-assortative manner. Specifically, Corollary˜1 establishes that any optimal segmentation satisfies a weak form of monotonicity, which requires that, if a buyer consumes at a given price, there must exist a buyer with a greater willingness to pay who consumes at a higher price. This notion of progressive pricing is mild, since it still allows for instances in which some buyer with higher willingness to pay consumes at a price lower than some buyer with lower willingness to pay. We show in Proposition˜2 how a strengthening of redistributive concerns [^4] As shown below, this corresponds mathematically to the welfare function satisfying a strengthened notion of supermodularity (Definition 6). eliminates these instances, leading to a stronger notion of monotonicity: if a buyer consumes at a given price, then all buyers with a (weakly) greater willingness to pay consume at a higher price. Furthermore, we show that optimal strongly monotone segmentations are obtained by following a simple greedy algorithm that iteratively maximizes the welfare of every type starting from the lowest.
Second, while redistribution does not induce an equity-efficiency trade-off (Lemma˜1 shows that optimal redistributive segmentations are efficient), we show that it may be in conflict with consumer surplus maximization. The monotonic structure of redistributive segmentations implies separation of buyers with low and high willingness to pay into different segments, possibly allowing the seller to price discriminate more effectively than under consumer surplus maximizing segmentations. As a result, the seller might obtain a strictly positive share of the surplus generated by segmenting the market. In such cases, we say that the seller enjoys a redistributive rent. Redistributive rents arise especially when optimal segmentations satisfy the strong form of monotonicity described in the previous paragraph. Proposition˜3 characterizes the markets for which strongly redistributive preferences necessarily induce a redistributive rent for the seller. When rents are not necessary, the optimal segmentation is remarkably simple. It only generates two segments: one where the lowest possible price is charged, pooling all types that would not have bought the good in the unsegmented market; and one segment where the uniform monopoly price is charged.
Our analysis considers the problem faced by a designer with redistributive concerns who can flexibly segment the aggregate market. In practice, however, market segmentation is primarily conducted by private actors whose objectives are not necessarily aligned with welfare maximization. For example, online platforms can focus on maximizing the number of transactions (Romanyuk2019), while data intermediaries aim to maximize profits (Yang2022). In the last section of the paper, we therefore investigate under what conditions redistributive segmentations can be implemented by a regulator who can only imperfectly observe the segmentation chosen by an agent, assumed here to be the seller. [^5] The result would still hold if the agent had a different objective function, provided it is aligned with efficiency. For example, a platform remunerated by a flat fee on transactions would seek to maximize participation, and thus segment efficiently. We introduce a notion of price-based implementation: a segmentation is implementable through price-based regulation if there exists no alternative segmentation that strictly increases the seller’s profit while inducing the same distribution of prices. [^6] This notion of implementation is equivalent to linliu’s (linliu) notion of credibility. We further discuss the connection to that contribution in Section 4, Footnote 35. This notion of implementation is motivated by the fact that although regulators are likely to possess less information than data intermediaries, they can plausibly observe the distribution of charged prices. [^7] Many websites, such as Camelcamelcamel, Keepa, Price Before, and Honey, already specialize in tracking the evolution, frequency and extent of price fluctuations for products sold on Amazon. In the context of regulating online platforms, for instance, it only involves monitoring transaction data and imposing penalties based on the divergence of the empirical price distribution from the desired target distribution. In Proposition˜4, we show that any optimal redistributive segmentation is implementable through price-based regulation. [^8] In fact, Proposition 4 states a more general result of independent interest: any efficient segmentation can be implemented via price-based regulation.
Related Literature
Our paper contributes to the analysis of the welfare effects of third-degree price discrimination, following pigou and robinson. Much of this literature focuses on studying how given market segmentations affect efficiency, consumer surplus, total output, and prices (schmalensee; varian; Aguirre2010Monopoly; Cowan2016Welfare; rhodeszhou).
legrand is the first to study theoretically the use of third-degree price discrimination as a redistributive device for public firms. His analysis is expanded by steinberg to the case where there is competition between for-profit firms and a (distributionally minded) non-profit provider.
bbm15 pioneered an information design approach to third-degree price discrimination. [^9] See kamenicareview and Bergemann2019 for comprehensive reviews of the Bayesian persuasion and information design literature. They show that any pair of consumer surplus and producer profit is attainable through some segmentation if the seller’s profit is at least the uniform-pricing level, consumer surplus is non-negative, and total surplus does not exceed the efficient level. hagpanah_siegel_aer and terstiege2023 extend this analysis to multi-product environments, elliott2024Market and bergemann2024alignment study competitive markets, and Salas2025 examines how optimal segmentation is shaped by endogenous buyer participation. We contribute to this literature by studying the impact of market segmentation on consumer welfare beyond the standard of consumer surplus. In particular, we study redistributive welfare functions and show that prioritizing low-type consumers does not hinder efficiency but may not maximize total consumer surplus.
A burgeoning literature rooted in the seminal contribution of weitzman1977 shows how markets (dworczak21), allocation mechanisms (Condorelli2013; Reuter2020; akbarpour2024vaccines; akbarpour2024redistributive; kang2023; kang2024a; kang2024b), or commodity taxation (pai2022taxing; Ahlvik2025; Doligalski2025) can be optimally designed to mitigate economic inequality.101010dworczak2024inequality provides an overview of recent contributions to inequality-aware market design. Our paper contributes to this literature by showing how market segmentation (or, equivalently, information) can serve as a redistributive device.111111arya2022rethinking study redistribution through a different notion of market segmentation, namely the non-tradability of essential and non-essential goods in a general equilibrium framework.
Relatedly, doval_smolin_22 analyze how information provision affects the welfare of heterogeneous agents. They introduce the Bayes welfare set, defined as the set of agents’ welfare profiles that can be achieved under some information policy. Their framework applies to a broad class of persuasion environments, including the monopolistic market we study. They characterize this set geometrically and show that its Pareto frontier is generated by the solutions to a family of Bayesian persuasion problems in which a weighted-utilitarian social planner plays the role of the designer. This delivers a complete description of the welfare effects that any segmentation can have on consumers of different types.
Rather than characterizing which welfare profiles are feasible, we study how to optimally distribute consumer welfare through information design. To do so, we specify a class of redistributive welfare functions—those that are supermodular in the types and the prices—and fully characterize all segmentations that are optimal for such objectives, for any distribution of consumer types. This characterization implies the use of monotone information structures as optimal policies—structures that are often viewed as especially attractive in the persuasion literature because of their simplicity (see, e.g., Ivanov2021; Mensch2021; Nikzad2021; Kolotilin2022; Onuchic2023; Kolotilin2024). This literature often obtains monotone structures as optimal in the case where payoffs are linear in the state (that is, only the posterior mean matters). [^12] Notable exceptions are, e.g., Mensch2021 and Nikzad2021. This is not the case in the environment we consider—both the sender’s and receiver’s payoffs are non-linear. We thus develop a new method inspired by the literature on stochastic orders (cf. Proposition˜1) to prove our characterization. [^13] We discuss the connection between our approach and the literature on stochastic orders and optimal transport in greater detail in Section 3.5.2.
The distributive impacts of price discrimination have also been investigated empirically. Dube2023 conduct a randomized controlled field experiment to study the welfare implications of personalized pricing. While they find that personalized pricing reduces total consumer surplus, aggregate consumer welfare can increase under inequality-averse welfare criteria, indicating that buyers with low willingness to pay benefit the most from the implemented price discrimination. Similarly, buchholz2020value examine price discrimination on a two-sided platform for cab rides and find that, although aggregate consumer surplus falls, more price-sensitive buyers gain from personalized pricing.
2. Model
Primitives
A monopolistic seller (she) supplies a good to a unit mass of buyers. Buyers have a unit demand and differ in their willingness to pay for the good, $\theta$, assumed to belong to the finite set $\Theta=\{\theta_{1},\dots,\theta_{K}\}\subset\mathbb{R}_{+}$, where $\theta_{1}<\dots<\theta_{K}$. We refer to the willingness to pay as the buyer’s type and use the two terms interchangeably. A buyer willing to pay $\theta$ consumes at a price $p\in\mathbb{R}_{+}$ if and only if $\theta\geq p$. We assume that the seller’s marginal cost of production is zero, so the gains from trade are positive. [^14] This assumption is without loss of generality. If the seller’s marginal cost were equal to $c>0$, then one could redefine the buyer type as $\tilde{\theta}=\max\{0,\theta-c\}$ for all $\theta\in\Theta$. The aggregate market (in short the market) $\mu\in\Delta(\Theta)$ specifies the initial distribution of types.
Market segmentation
A market segmentation is a division of the population of buyers into different market segments indexed by $s\in S$, where $S$ is a sufficiently large finite set. Formally, a segmentation corresponds to a probability distribution $\sigma\in\Delta(\Theta\times S)$ describing the mass $\sigma(\theta,s)$ of type $\theta$ buyers assigned to segment $s$, and satisfying the following market consistency constraint:
$\forall\theta\in\Theta,\quad\sum_{s\in S}\sigma(\theta,s)=\mu(\theta).$ ($\mathrm{C}_{\mu}$)
Condition $\mathrm{C}_{\mu}$ ensures that the mass of type $\theta$ buyers assigned to all segments coincides with the mass of those types in the market. [^15] We can equivalently think of any pair $(\sigma,S)$ as a Blackwell experiment which, for any prior distribution $\mu$ over the state space $\Theta$, describes the joint distribution of signals $s\in S$ and states of the world $\theta\in\Theta$, and satisfies the Bayes-plausibility constraint $\mathrm{C}_{\mu}$ (Aumann1995; kamenica_gentzkow11). According to that interpretation, each $\sigma$ represents a way of revealing information about the buyers’ types to the seller.
Seller’s pricing
When faced with a market segmentation, the seller sets prices to maximize her profit in each segment. That is, for any $\mu\in\Delta(\Theta)$, any $\sigma$ satisfying $\mathrm{C}_{\mu}$, and any $s\in S$ such that $\sigma(\theta,s)>0$ for some $\theta\in\Theta$, the seller solves
$\max_{p\in P}\;p\sum_{\theta\geq p}\sigma(\theta,s),$
where the set of prices is given by $P=\Theta$. Indeed, there is no loss of generality in assuming that the seller charges prices in $\Theta$ since, for any $k\in\{1,\dots,K\}$, charging a price falling strictly between two consecutive types $\theta_{k}$ and $\theta_{k+1}$ always yields a lower profit to the seller than charging either $p=\theta_{k}$ or $p=\theta_{k+1}$. [^16] Similarly, charging any price lower than $\theta_{1}$ or greater than $\theta_{K}$ is never optimal.
Direct and obedient segmentations
It is without loss of generality to focus on direct and obedient market segmentations. [^17] Precisely, it is without loss given the objectives we define below. See Bergemann2019. In such segmentations, $S=P$, so that the segments are indexed by a recommended price, and the seller finds it optimal to follow this recommendation. That is:
$\forall p,q\in P,\quad p\sum_{\theta\geq p}\sigma(\theta,p)\geq q\sum_{\theta\geq q}\sigma(\theta,p).$ (Ob)
For any $\mu\in\Delta(\Theta)$, we denote the set of direct and obedient segmentations of $\mu$ as $\Sigma(\mu)$. We henceforth slightly abuse terminology by referring to any element of $\Sigma(\mu)$ as a segmentation of $\mu$. For any $\mu\in\Delta(\Theta)$ and any $\sigma\in\Sigma(\mu)$, we let $\sigma_{P}\in\Delta(P)$ be the distribution of recommended prices induced by $\sigma$, defined by: [^18] As will become clear after Lemma 1 is presented, in optimal segmentations the distribution of recommended prices will coincide with the distribution of realized transaction prices.
$\forall p\in P,\quad\sigma_{P}(p)=\sum_{\theta\in\Theta}\sigma(\theta,p).$
For any $\mu\in\Delta(\Theta)$, we let $p^{\star}_{\mu}\in P$ be the uniform optimal price, that is, the price the seller would optimally charge in the aggregate market: [^19] Whenever there are several uniform optimal prices for some market $\mu\in\Delta(\Theta)$, we let $p^{\star}_{\mu}$ be the lowest one. This tie-breaking rule is inconsequential for any of our results.
$\forall p\in\operatorname{supp}(\mu),\quad p^{\star}_{\mu}\sum_{\theta\geq p^{\star}_{\mu}}\mu(\theta)\geq p\sum_{\theta\geq p}\mu(\theta).$
Let $\mu\in\Delta(\Theta)$. An example of direct segmentation of $\mu$ is illustrated in Figure˜1.
[…]
Figure 1: A direct segmentation with two segments.
The initial mass of consumers, located in the row $p^{\star}_{\mu}$ in the grid $\Theta\times P$, is divided between two segments $p^{\prime}$ and $p^{\prime\prime}$ vertically, so $\mathrm{C}_{\mu}$ is satisfied. Furthermore, Ob requires that the mass moved at each point be such that $p^{\prime}$ and $p^{\prime\prime}$ are, respectively, profit-maximizing prices in segments $p^{\prime}$ and $p^{\prime\prime}$.
Redistributive welfare functions
We consider a designer (he) who can segment the market flexibly. His preferences over market outcomes are represented by a welfare function $w\colon\Theta\times P\to\mathbb{R}$. For any welfare function $w$ and segmentation $\sigma\in\Sigma(\mu)$, the aggregate welfare is defined by
$\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma(\theta,p).$
The value $w(\theta,p)$ thus represents the individual contribution to the aggregate welfare when a buyer of type $\theta$ is assigned to a segment where the seller charges the price $p$.
We restrict attention to welfare functions satisfying three conditions. First, the value from assigning a buyer to a segment where he does not consume is zero, and the value from assigning a buyer to a segment where he consumes is non-negative. That is, $w(\theta,p)=0$ if $\theta<p$, and $w(\theta,p)\geq 0$ if $\theta\geq p$. Second, assigning any buyer to a segment where the seller charges a lower price always improves welfare. That is, $p\mapsto w(\theta_{k},p)$ is a non-increasing function on $\{\theta_{1},\dots,\theta_{k}\}$ for all $k\in\{1,\dots,K\}$. The two preceding conditions capture the fact that social welfare aligns with consumer surplus. [^20] It is straightforward to verify that the buyer-utilitarian welfare function, given by $w(\theta,p)=(\theta-p)\mathds{1}_{\theta\geq p}$ for all $(\theta,p)\in\Theta\times P$, satisfies these two minimal conditions. The third condition captures the designer’s redistributive concerns. We require that any welfare function be supermodular conditional on buyers consuming. That is, for any $(\theta,p),(\theta^{\prime},p^{\prime})\in\Theta\times P$ such that $\theta^{\prime}>\theta\geq p^{\prime}>p$,
$w(\theta,p)-w(\theta,p^{\prime})\geq w(\theta^{\prime},p)-w(\theta^{\prime},p^{\prime}).$ (R)
Condition R ensures that any price discount from $p^{\prime}$ to $p$ increases welfare more when it is granted to a low-type $\theta$ rather than to a high-type $\theta^{\prime}$.
We say that a welfare function is redistributive if it satisfies the three conditions above, and denote by $\mathcal{R}$ the class of those welfare functions. Furthermore, we say that $w$ is strictly redistributive if it satisfies the condition R strictly on all $\Theta\times P$, and is such that $p\mapsto w(\theta_{k},p)$ is strictly decreasing on $\{\theta_{1},\dots,\theta_{k}\}$ for all $k\in\{1,\dots,K\}$. We denote the class of strictly redistributive welfare functions by $\overline{\mathcal{R}}$.
In Examples˜1 and 2 below, we highlight that the class of redistributive welfare functions encompasses two common specifications.
Example 1 (Decreasing Pareto weights).
Let $\lambda\colon\Theta\to\mathbb{R}$ be such that $\lambda(\theta)\geq 0$ for any $\theta\in\Theta$. The welfare function defined by
$\forall(\theta,p)\in\Theta\times P,\quad w(\theta,p)=\lambda(\theta)(\theta-p)\mathds{1}_{\theta\geq p},$ (1)
belongs to $\mathcal{R}$ whenever $\theta\mapsto\lambda(\theta)$ is non-increasing, and belongs to $\overline{\mathcal{R}}$ whenever $\theta\mapsto\lambda(\theta)$ is strictly decreasing. The class of welfare functions given by 1 for some decreasing function $\lambda$ corresponds to the family of weighted buyer-utilitarian welfare functions with decreasing Pareto weights. Under such welfare functions, the value to the designer of an additional unit of surplus is lower for high-types than for low-types. Decreasing Pareto weights can endogenously arise under a utilitarian objective if buyers with a low willingness to pay have a higher marginal utility of money than those with a high willingness to pay (Condorelli2013; dworczak21), or if buyers are identified by unobserved welfare weights that negatively correlate with their willingness to pay (akbarpour2024redistributive).
Example 2 (Increasing concave transformation).
Let $u\colon\mathbb{R}\to\mathbb{R}$ be an increasing function such that $u(0)=0$. The welfare function defined by
$\forall(\theta,p)\in\Theta\times P,\quad w(\theta,p)=u\bigl((\theta-p)\mathds{1}_{\theta\geq p}\bigr),$ (2)
belongs to $\mathcal{R}$ if $u$ is concave, and belongs to $\overline{\mathcal{R}}$ if $u$ is strictly concave. Increasing concave transformations of surplus correspond to the class of welfare functions usually considered in optimal taxation (mirrlees1971exploration) and insurance (Rothschild1976). Concavity reflects equity concerns, as it implies diminishing marginal returns to welfare from additional units of surplus.
Note also that any welfare function which is a combination of Examples˜1 and 2, given by
$\forall(\theta,p)\in\Theta\times P,\quad w(\theta,p)=\lambda(\theta)u\bigl((\theta-p)\mathds{1}_{\theta\geq p}\bigr),$
belongs to $\mathcal{R}$ whenever $\theta\mapsto\lambda(\theta)$ is sufficiently decreasing, or $u$ is sufficiently concave, or both.
Segmentation design
For any given welfare function $w\in\mathcal{R}$ and an aggregate market $\mu\in\Delta(\Theta)$, the designer’s optimization program consists of choosing a segmentation of $\mu$ to maximize aggregate welfare:
$\max_{\sigma\in\Sigma(\mu)}\;\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma(\theta,p).$ ($\mathrm{P}_{w,\mu}$)
3. Redistributive Segmentations and Welfare
We start by proving a preliminary result that simplifies the analysis: it is without loss of generality to restrict the search for optimal redistributive segmentations to the efficient ones. We then characterize the optimal redistributive segmentations. Finally, we discuss the economic implications of our characterization. First, we show that optimal redistributive segmentations must exhibit a monotonic structure that induces the seller to price progressively. Second, we show that when the redistributive concerns are strong, optimal redistributive segmentations can lead the seller to obtain a strictly positive share of the surplus.
3.1.  Efficiency
A segmentation is efficient if it leads the seller to serve all the buyers. Let $\Omega=\bigl\{(\theta,p)\in\Theta\times P\ |\ \theta\geq p\bigr\}$. For any $\mu\in\Delta(\Theta)$, the set of efficient segmentations of $\mu$ is defined by
$\Sigma^{\mathrm{eff}}(\mu)=\bigl\{\sigma\in\Sigma(\mu)\ |\ \operatorname{supp}(\sigma)\subseteq\Omega\bigr\}.$
Lemma 1 (Efficiency).
For any $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$, there exists an optimal segmentation for the problem $\mathrm{P}_{w,\mu}$ that belongs to $\Sigma^{\mathrm{eff}}(\mu)$.
Proof.
Let $\mu\in\Delta(\Theta)$ and consider $\sigma\in\Sigma(\mu)$ with $\operatorname{supp}(\sigma)\not\subset\Omega$. Therefore, there must exist $(\theta,p)\in\operatorname{supp}(\sigma)$ such that $\theta<p$. Consider reallocating the entire mass of buyers $\sigma(\theta,p)$ to $(\theta,\theta)$. This reallocation violates neither $\mathrm{C}_{\mu}$ nor Ob. The total mass of buyers of type $\theta$ is still equal to $\mu(\theta)$ so $\mathrm{C}_{\mu}$ is preserved. Furthermore, this reallocation relaxes Ob in both segments $\theta$ and $p$. Finally, since $w(\theta,p)=0\leq w(\theta,\theta)$, the aggregate welfare must increase. In fact, aggregate welfare weakly increases without hurting the seller or the buyers—a (weak) Pareto improvement. ∎
Lemma˜1 establishes that segmenting a market in a redistributive way never hinders efficiency. The intuition behind that result is similar to hagpanah_siegel_jpe’s (hagpanah_siegel_jpe) main result: any inefficient segmentation is Pareto-dominated (with respect to the seller’s profit and consumer surplus) by some efficient one. For our class of welfare functions, such Pareto improvements always increase aggregate welfare, since $w(\theta,\,\cdot\,)$ is a decreasing function for each $\theta\in\Theta$.
Operationally, Lemma˜1 implies that we can restate the segmentation design problem $\mathrm{P}_{w,\mu}$ as follows:
$\max_{\sigma\in\Sigma^{\mathrm{eff}}(\mu)}\;\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma(\theta,p).$ ($\mathrm{P}^{\star}_{w,\mu}$)
For any $w\in\mathcal{R}$ and $\mu\in\Delta(\Theta)$, we say that $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal if it solves the optimization program $\mathrm{P}^{\star}_{w,\mu}$. We call optimal redistributive segmentations all segmentations that are $\mathrm{P}^{\star}_{w,\mu}$-optimal for some $\mu\in\Delta(\Theta)$ and $w\in\mathcal{R}$.
3.2.  Characterization of Optimal Segmentations
The characterization of $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentations, stated below as Theorem˜1, is given in terms of the conditions described in Definition˜1 (see below for a detailed explanation).
Definition 1 (Redistributive Saturation).
Let $\mu\in\Delta(\Theta)$. A segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is redistributively saturated (saturated for short) if the following conditions hold:
  1. (a)
    In any segment $p\in\operatorname{supp}(\sigma_{P})$ such that $p<\max\operatorname{supp}(\sigma_{P})$, the seller is indifferent between charging $p$ and some other price $q>p$;
  2. (b)
    Let $p\in\operatorname{supp}(\sigma_{P})$ and $\theta\in\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)$. Then in every segment $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}\leq\theta$ the seller is indifferent between charging $p^{\prime}$ and $\theta$.
An efficient segmentation is saturated if (a) in every segment, except maybe for the one with the highest price, the seller faces more than one optimal price, and (b) if some consumers of type $\theta$ belong to a segment with price $p$, then consumers of type $\theta$ must be sufficiently represented in higher segments so that $\theta$ is an optimal price in those segments.
Theorem 1 (Optimal redistributive segmentations).
For any $\mu\in\Delta(\Theta)$, the two following claims are true:
  1. (i)
    If $w\in\mathcal{R}$, then there exists a $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation which is saturated. Furthermore, if $w\in\overline{\mathcal{R}}$, then every $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is saturated.
  2. (ii)
    Conversely, if $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is saturated, then there exists $w\in\mathcal{R}$ such that $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal.
The above result establishes that the set of saturated segmentations and optimally redistributive ones are (almost) equal. Claim (i) states that an optimal segmentation must be saturated if the welfare function is strictly redistributive ($w\in\overline{\mathcal{R}}$). If $w\in\mathcal{R}$ is not strictly redistributive, then there may be indifferences that make some non-saturated segmentations also optimal, but a saturated optimal one always exists. Claim (ii) states that, conversely, any saturated segmentation is optimal for some redistributive objective. As a result, any property that is satisfied by saturated segmentations is also satisfied by optimal redistributive segmentations. We leverage this result to draw economic implications of optimal redistributive segmentations in Sections˜3.3 and 3.4; Section˜3.2.2 outlines its proof. [^21] From now on, with slight abuse of terminology, we use the terms optimally redistributive segmentation and saturated segmentation interchangeably.
3.2.1. Intuitive Explanation of Theorem˜1
Theorem˜1 is best understood in light of what saturated segmentations achieve. Saturated segmentations are those for which no reallocation of buyers between segments could increase the expectation of every redistributive welfare function without leading to a violation of the seller’s obedience constraints. At this stage, two questions are in order: (1) Which type of reallocation increases the expectation of every redistributive welfare function? (2) When is it not possible to perform such reallocations without violating the seller’s obedience constraints?
Proposition˜1, formally stated in Section˜3.2.2, answers question (1) by identifying the two marginal reallocations of buyers that increase any redistributive welfare function. First, downward buyer transfers (Definition˜3) move buyers to segments with lower prices. Second, redistributive buyer transfers (Definition˜4) swap a given mass of low-type buyers in a high-price segment for an equal mass of higher-type buyers in a lower-price segment. Figure˜2 illustrates both transfers on an efficient segmentation.
[…]
Figure 2: The shaded region represents the set $\Omega$. A downward transfer involving a mass $\delta$ of buyers (in blue) and a redistributive transfer involving a mass $\varepsilon$ of buyers (in orange).
Downward transfers are desirable under redistributive welfare functions as they decrease the price charged to some buyers. Redistributive transfers are also desirable under redistributive welfare functions due to condition R. The main contribution of Proposition˜1 is to establish the converse: any buyer reallocation that increases the expected value of all redistributive welfare functions must necessarily consist of some (finite) combination of downward and redistributive buyer transfers. This result provides a complete characterization of welfare-improving buyer reallocations under redistributive objectives.
While downward and redistributive buyer transfers are the key to making segmentations more redistributive, these transfers interfere with the seller’s obedience constraints. Reallocating buyers through downward or redistributive transfers can lead the seller to deviate to a higher price in one or more segments. Lemmas˜2 and 3, also formally stated in Section˜3.2.2, answer question (2). They show that saturation is precisely the condition that ensures that no downward or redistributive transfer (or a combination thereof) can be performed without causing the seller to deviate from the recommended prices.
To get a sense of why, let us first investigate the conditions under which a particular downward transfer is infeasible for some segmentation. Consider transferring some mass of buyers of type $\tilde{\theta}$ from a higher-price segment $p^{\prime\prime}$ to a lower-price segment $p^{\prime}$, as illustrated in Figure˜3(a).
[…]
(a) Downward transfer.
[…]
(b) Redistributive transfer.
Figure 3: Points within the orange boxes represent $(\theta,p)$ for which Ob tightens after a transfer.
This transfer affects the seller’s obedience constraints in both segments. In the lower segment $p^{\prime}$, adding these buyers increases demand at prices between $p^{\prime}$ and $\tilde{\theta}$, tightening constraints for these prices (shown in the orange box in the lower segment). Simultaneously, in the higher segment $p^{\prime\prime}$, removing buyers decreases demand at prices between $p^{\prime\prime}$ and $\tilde{\theta}$, tightening constraints for prices above $\tilde{\theta}$ (represented in the orange box in the upper segment). If at least one of the obedience constraints is already binding at those prices in either segment, even a small transfer will violate the seller’s obedience constraints, making the downward transfer infeasible.
The reasoning is similar for redistributive transfers. Consider swapping some mass of buyers between segments as shown in Figure˜3(b): buyers of type $\theta^{\prime}$ move from segment $p^{\prime\prime}$ to $p^{\prime}$, while buyers of type $\theta^{\prime\prime}$ move from $p^{\prime}$ to $p^{\prime\prime}$. This transfer does not tighten any of the obedience constraints in segment $p^{\prime}$. Indeed, it decreases the demand at all prices between $\theta^{\prime}$ and $\theta^{\prime\prime}$, and does not affect the demand at any other price. However, in the higher segment $p^{\prime\prime}$, it increases demand at all these intermediate prices (highlighted by the orange box in Figure˜3(b)), potentially violating obedience constraints in that segment.
Saturated segmentations are precisely those for which no downward or redistributive transfers remain feasible. Condition (ii) ensures that no buyer of type $\theta$ belonging to segment $p$ can be moved upwards, as the obedience constraint for price $\theta$ binds in every upper segment they could be moved to. [^22] Since we are restricting attention to efficient segmentations, this corresponds to segments $p^{\prime}$ such that $p<p^{\prime}\leq\theta$. This blocks all redistributive transfers, which necessarily involve moving some buyers upward. Condition (ii) also blocks some downward transfers: one cannot transfer some type $\theta$ downwards to a segment that includes any higher type $\theta^{\prime}>\theta$, since this would lead to a violation of obedience in the origin segment. Transferring $\theta$ downwards to a segment without any higher types, however, could in principle be possible. Condition (i) blocks those remaining downward transfers by ensuring that even if such transfer does not lead to a violation of obedience in the origin segment, it does in the destination one.
3.2.2. The Proof of Theorem˜1
The proof of Theorem˜1 is based on a key result of independent interest. We articulate a sense in which one segmentation is more redistributive than another by introducing a partial order on the space of efficient segmentations (Definition˜2). Under this order, a segmentation is more redistributive than another if the former generates higher aggregate welfare than the latter for every redistributive welfare function. Proposition˜1 establishes that this occurs precisely when the more redistributive segmentation can be obtained through a sequence of downward and redistributive buyer transfers applied to the less redistributive one. [^23] This provides an answer to a practical question: Say one observes some market segmentation. How can it be modified to make it more redistributive? We then demonstrate that optimal redistributive segmentations are exactly those for which no segmentation that is more redistributive exists (Lemma˜2) and that this condition is equivalent to being saturated (Lemma˜3).
Proof of Theorem˜1.
We first introduce a redistributive order on the space of efficient segmentations: [^24] See Section 3.5.2 for the detailed connection between our redistributive order and existing stochastic orders, in particular, the supermodular one (shaked2007; Muller2013; Meyer2015).
Definition 2 (Redistributive order).
Let $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$. We say that $\sigma$ is more redistributive than $\sigma^{\prime}$, denoted $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$, if
$\forall w\in\mathcal{R},\quad\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma(\theta,p)\geq\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma^{\prime}(\theta,p).$
[…]
We call a buyer mass transfer (in short, a transfer) any $t\colon\Omega\to\mathbb{R}$ that satisfies $\sum_{p\leq\theta}t(\theta,p)=0$ for any $\theta\in\Theta$. Intuitively, transfers perturb a segmentation $\sigma$ by moving buyers between segments, with $t(\theta,p)$ denoting the mass of buyers of type $\theta$ that is added to or removed from a segment with price $p$. We now define two particular classes of transfers.
Definition 3 (Downward transfers).
A transfer $t$ is downward if there exist $\delta\geq 0$ and $\tilde{\theta},p^{\prime},p^{\prime\prime}\in\Theta$ such that $p^{\prime}<p^{\prime\prime}$, and
$t(\theta,p)=\left\{\begin{array}[]{ll}\delta&\text{if $(\theta,p)=(\tilde{\theta},p^{\prime})$}\\ -\delta&\text{if $(\theta,p)=(\tilde{\theta},p^{\prime\prime})$}\\ 0&\text{otherwise}\end{array}\right$.
Definition 4 (Redistributive transfers).
A transfer $t$ is redistributive if there exists $\varepsilon\geq 0$ and $(\theta^{\prime},p^{\prime\prime}),(\theta^{\prime\prime},p^{\prime})\in\Omega$ such that $\theta^{\prime}<\theta^{\prime\prime}$, $p^{\prime}<p^{\prime\prime}$, and
$t(\theta,p)=\left\{\begin{array}[]{ll}\varepsilon&\text{if $(\theta,p)\in\bigl\{(\theta^{\prime},p^{\prime}),(\theta^{\prime\prime},p^{\prime\prime})\bigr\}$}\\[5.0pt] -\varepsilon&\text{if $(\theta,p)\in\bigl\{(\theta^{\prime},p^{\prime\prime}),(\theta^{\prime\prime},p^{\prime})\bigr\}$}\\[5.0pt] 0&\text{otherwise}\end{array}\right$.
Proposition˜1 shows that performing a sequence of such transfers is not only sufficient to make a segmentation more redistributive (according to Definition˜2) but also necessary. Let $\mathcal{T}$ be the set of all downward and redistributive transfers.
Proposition 1.
Let $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$. The two following claims are equivalent:
  1. (i)
    $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$;
  2. (ii)
    There exists $n\in\mathbb{N}$, $n>0$, and $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that
    $\sigma=\sigma^{\prime}+\sum_{i=1}^{n}t_{i}.$
The proof of Proposition˜1, which can be found in Appendix˜A, relies on duality arguments akin to those employed by Muller2013 and Meyer2015. Specifically, we show that the space of redistributive welfare functions (up to an additive constant) is dual to the cone of downward and redistributive transfers.
[…]
Lemma 2.
For any $\mu\in\Delta(\Theta)$, the following two claims are true:
  1. (i)
[…]
  2. (ii)
[…]
Lemma 3.
[…]
$3.$  Monotonicity and Progressive Pricing
In this subsection, we show an important economic implication of Theorem˜1. Optimal redistributive segmentations exhibit a positive assortative structure. They assign low-type buyers to low-price segments and high-type buyers to high-price segments. Furthermore, this positive-assortative structure is strenghtened as redistributive concerns grow.
3.3.1. Weak Monotonicity
We start by introducing a weak notion of positive assortativity for segmentations:
Definition 5 (Weakly monotone segmentations).
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is weakly monotone if for any $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$,
$\max\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)\leq\max\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p^{\prime})\bigr).$
[…]
(a) Weak monotonicity.
[…]
(b) Strong monotonicity.
Figure 4: Monotone segmentations.
A segmentation is weakly monotone if, for any pair of segments, the highest type assigned to the lower-price segment is not strictly greater than the highest type assigned to the higher-price segment. Figure˜4(a) illustrates weakly monotone segmentations.
A direct implication of Theorem˜1 is that all optimal redistributive segmentations are weakly monotone. Indeed, by condition (b) in Definition˜1, for any saturated segmentation, if a type $\theta$ is in a segment $p\in\operatorname{supp}(\sigma_{P})$, then it must belong to any other segment $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}\leq\theta$, which implies weak monotonicity. [^26] We state the result for strictly redistributive welfare functions to clarify the exposition. If $w\in\mathcal{R}$, then for any $\mu\in\Delta(\Theta)$, there exists a $\mathrm{P}^{\star}_{w,\mu}$-optimal and weakly monotone segmentation of $\mu$.
Corollary 1.
For any $\mu\in\Delta(\Theta)$ and $w\in\overline{\mathcal{R}}$, every $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is weakly monotone.
Corollary˜1 establishes that redistributive segmentations induce a form of progressive pricing, whereby high-types are charged higher prices than low-types. In weakly monotone segmentations, for every buyer of a given type, there exists a buyer of a higher type that is charged a (weakly) higher price. This notion of progressivity is mild, since some buyers of a lower type might still pay higher prices than some buyers of higher type. This outcome arises in regions where the support of different segments overlap, i.e., when buyers are randomly assigned to segments.
[…]
(a) A saturated weakly monotone segmentation.
[…]
(b) A compensated redistributive transfer.
Figure 5: The left panel illustrates a saturated segmentation where some buyers $\theta_{k+1}$ consume at a lower price than some buyers $\theta_{k}$. Points in orange represent $(\theta,p)$ for which Ob binds.
3.3.2. Strengthening Monotonicity
Compensated redistributive transfers
Regions of random assignment are exactly where unresolved redistributive tensions remain in saturated segmentations. Consider the segmentation shown in Figure˜5(a): some buyers of type $\theta_{k}$ are placed in the higher segment to “make room” for some buyers of type $\theta_{k+1}$ in the lower segment. These $\theta_{k}$ buyers would be better off in the lower segment, but in a saturated segmentation this cannot be achieved by simple downward or redistributive transfers without violating Ob. In particular, any redistributive transfer between types $\theta_{k}$ and $\theta_{k+1}$ would induce the seller to deviate to price $\theta_{k+1}$ in the upper segment.
Yet such a reallocation of types $\theta_{k}$ to the lower segment can be feasibly implemented through an appropriate sequence of transfers. Consider the construction illustrated in Figure˜5(b): first, a redistributive transfer swaps a mass $\varepsilon$ of types $\theta_{k}$ and $\theta_{k+1}$; then an upward transfer moves a mass $\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\varepsilon$ of higher-type buyers $\theta_{\ell}$ to a segment where they are charged exactly their willingness to pay. Removing this mass of type $\theta_{\ell}$ exactly offsets the effect of adding an $\varepsilon$ mass of types $\theta_{k+1}$ to segment $\theta_{k}$, preserving obedience. Moreover, because the transferred $\theta_{\ell}$ buyers are the highest types in segment $\theta_{k}$ and are moved to $(\theta_{\ell},\theta_{\ell})$, Ob remains satisfied. Thus, whenever a segmentation contains multiple types at the overlap between segments, one can construct such a feasible combination of redistributive and upward transfers. We refer to this construction as compensated redistributive transfers.
Strongly redistributive welfare functions
Compensated redistributive transfers may be desirable under welfare functions that reflect strong redistributive concerns. That is, welfare functions under which the welfare gain from lowering the price paid by a buyer $\theta_{k}$ offsets not only the welfare loss from an equivalent price increase for a buyer $\theta_{k+1}$ but also potential welfare losses borne by other higher types. Specifically, the compensated redistributive transfer illustrated in Figure˜5 is beneficial whenever the net welfare gain from the redistributive transfer between types $\theta_{k}$ and $\theta_{k+1}$ exceeds by a factor of $\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}$ the welfare loss from raising the price paid by a buyer of type $\theta_{\ell}$ from $\theta_{k}$ to $\theta_{\ell}$.
We now define a class of welfare functions for which every compensated redistributive transfer is desirable:
Definition 6 (Strongly redistributive welfare functions).
A welfare function $w\in\overline{\mathcal{R}}$ is strongly redistributive if for any $p,\theta_{k},\theta_{\ell}\in\Theta$ such that $p<\theta_{k}<\theta_{\ell}$,
[…]
Strong monotonicity
Compensated redistributive transfers render any segmentation more positive-assortative by removing cases where a higher-type buyer pays less than a lower-type buyer. [^27] Note that performing a compensated redistributive transfer on a saturated segmentation also increases the seller’s profit, since the upward transfer increases the price being charged to $\theta_{\ell}$ while the redistributive transfer is profit-neutral. This aspect of compensated redistributive transfers is further examined in Section 3.4. As redistributive concerns become stronger, more compensated redistributive transfers should be performed. This process reaches a limit once no further compensated redistributive transfers can be implemented, yielding what we call strongly monotone segmentations:
Definition 7 (Strongly monotone segmentations).
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is strongly monotone if for any $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$,
$\max\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)\leq p^{\prime}.$
A segmentation is strongly monotone if buyers are assigned to segments deterministically, except possibly for types lying at the boundary between two consecutive segments. [^28] Our definition of weak monotonicity implies that segment supports are weak-set-ordered (Che2021). Strong monotonicity, by contrast, induces an ordering of segment supports stronger than the strong set order (topkis). Since the intersection of any two segment supports contains at most one element, strongly monotone segmentations correspond to a slightly weaker notion than monotone partitional information structures, which partition the state space into convex sets (Ivanov2021; Mensch2021; Kolotilin2022; Onuchic2023; Kolotilin2024). Strongly monotone segmentations, illustrated in Figure˜4(b), thus represent a pure form of progressive pricing: for any type, every higher type is charged a (weakly) higher price.
Optimality of strong monotonicity
Our next result shows that strongly redistributive welfare functions are the only ones for which every optimal segmentation is strongly monotone.
Proposition 2.
Any $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is strongly monotone for every $\mu\in\Delta(\Theta)$ if and only if $w\in\overline{\mathcal{R}}$ is strongly redistributive.
The proof of Proposition˜2 appears in Appendix˜D. We further show in the following lemma (the proof of which can be found in Appendix˜E) that, for any given aggregate market, there is exactly one strongly monotone segmentation of that market that is saturated. This particular segmentation therefore represents the uniquely optimal market segmentation for any strongly redistributive welfare function.
Lemma 4.
For any $\mu\in\Delta(\Theta)$, there exists a unique strongly monotone segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ that is saturated.
We now come back to Examples˜1 and 2. We show that the weighted consumer surplus welfare function is strongly redistributive if and only if Pareto weights decrease sufficiently fast in willingess to pay. [^29] This shows incidentally that the set of strongly redistributive welfare functions is non-empty. We also show that concave increasing transformations of consumer surplus are never strongly redistributive.
Example 0 (Decreasing Pareto weights, cont’d).
Let $\lambda\colon\Theta\to\mathbb{R}$ be a strictly decreasing function such that $\lambda(\theta)>0$ for all $\theta\in\Theta$. Define $w$ such that, for any $(\theta,p)\in\Omega$, $w(\theta,p)=\lambda(\theta)(\theta-p)$. By rearranging terms in the inequality 3, we can see that $w$ is strongly redistributive if and only if for any $\theta_{k},\theta_{\ell}\in\Theta$ such that $2\leq k\leq K-1$ and $k<\ell\leq K$,
$\frac{\lambda(\theta_{k})-\lambda(\theta_{k+1})}{\lambda(\theta_{\ell})}\geq\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\frac{\theta_{\ell}-\theta_{k}}{\theta_{k}-\theta_{k-1}}.$ (4)
Inequality 4 imposes a condition on the rate of decrease of $\theta\mapsto\lambda(\theta)$. In this case, we can reinterpret Proposition˜2 as follows: every optimal segmentation is strongly monotone if and only if Pareto weights decrease at a sufficiently fast rate. For instance, weights of the form $\lambda(\theta)=\theta^{-\gamma}$ or $\lambda(\theta)=\exp(-\gamma\theta)$ are strongly redistributive if $\gamma$ exceeds some threshold.
Example 1 (Increasing concave transformations, cont’d).
Let $u:\mathbb{R}\to\mathbb{R}$ be strictly increasing and strictly concave. Define $w$ such that, for any $(\theta,p)\in\Omega$, $w(\theta,p)=u(\theta-p)$. In this case, one can show that condition 3 can never be satisfied. In contrast to Example˜1, when the redistributive concerns stem from a concave transformation of buyers’ surplus, the welfare gain associated with redistributive transfers can never outweigh the welfare loss associated with burning the surplus of some high type. In this case, Proposition˜2 implies that there always exists a market $\mu\in\Delta(\Theta)$ such that no optimal segmentation is strongly monotone.
The greedy redistributive algorithm
The simple structure of strongly monotone segmentations and their uniqueness pave the way for a constructive approach. We describe a procedure, called the greedy redistributive algorithm, that outputs, for any market $\mu$, the unique strongly monotone segmentation of $\mu$ that is saturated. We explain in words the functioning of the algorithm and LABEL:secap:greedy contains its pseudo-code.
Strongly monotone segmentations can be seen as the result of lexicographic optimization: it benefits each type sequentially as much as possible, from the smallest to the largest. The greedy redistributive algorithm follows this logic and sequentially assigns types into segments. It starts by creating a lower segment with all buyers of type $\theta_{1}$ and then proceeds to include types $\theta_{2}$ in that segment. If all buyers $\theta_{2}$ can be included while keeping the price in the segment at $\theta_{1}$, the algorithm does so and proceeds to include types $\theta_{3}$. At some point in this process, the algorithm reaches a point such that including all buyers of a given type $\theta_{k}$ would violate Ob in that segment. When the algorithm reaches this point, it fits as many types $\theta_{k}$ as possible in the segment so that Ob becomes binding for some price and moves on to creating a new segment with the remaining mass of types $\theta_{k}$. This inclusion process is then repeated in the newly created segment starting at type $\theta_{k+1}$, and then proceeds sequentially as previously explained. The algorithm ends when every buyer has been assigned to some segment.
3.4.  Redistributive Rents
We now study how redistributive segmentations distribute surplus between the seller and the buyers. We show that if the welfare function is strongly redistributive, then optimal segmentation might allocate some of the created surplus to the seller. Hence, although the welfare function does not take into consideration the seller’s profit, redistributive segmentations may not maximize total consumer surplus.
For any $\mu\in\Delta(\Theta)$, we say that a segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ gives a rent to the seller if the profit under $\sigma$ is strictly higher than the profit under uniform pricing: that is, if
$\sum_{p\in P}p\sum_{\theta\geq p}\sigma(\theta,p)>p^{\star}_{\mu}\sum_{\theta\geq p^{\star}_{\mu}}\mu(\theta).$
The following joint probability distribution over types and prices will prove to be of particular interest. For any market $\mu\in\Delta(\Theta)$, we define $\sigma^{\star}_{\mu}\in\Delta(\Theta\times P)$ as follows:
[…]
Figure 6: The segmentation $\sigma^{\star}_{\mu}$ for a uniform price $p^{\star}_{\mu}$.
By construction, $\sigma^{\star}_{\mu}$ satisfies $\mathrm{C}_{\mu}$ and $\operatorname{supp}(\sigma^{\star}_{\mu})\subset\Omega$. Therefore, if $\sigma^{\star}_{\mu}$ satisfies Ob then it is a segmentation belonging to $\Sigma^{\mathrm{eff}}(\mu)$. In this case, we refer to $\sigma^{\star}_{\mu}$ as being feasible. Note that, if $\sigma^{\star}_{\mu}$ is feasible, then it is the unique saturated and strongly monotone segmentation of [^30] Note that $\sigma^{\star}_{\mu}$ is constructed such that, if it is feasible, the seller is indifferent between charging $\theta_{1}$ and $p^{\star}_{\mu}$ in segment $\theta_{1}$. As such, $\sigma^{\star}_{\mu}$ satisfies conditions (a) and (b) of Definition 1 if it is feasible, implying that it is saturated. Intuitively, the feasibility of $\sigma^{\star}_{\mu}$ is therefore demanding as it requires a large mass of type $p^{\star}_{\mu}$ such that $p^{\star}_{\mu}$ is an optimal price in both segments. $\mu$. Furthermore, $\sigma^{\star}_{\mu}$ has a very simple structure. It assigns all types lower than $p^{\star}_{\mu}$ to a “discount” segment where the lowest possible price $\theta_{1}$ is charged, and all the remaining types to a residual segment where the uniform price $p^{\star}_{\mu}$ is charged. Figure˜6 illustrates $\sigma^{\star}_{\mu}$.
Proposition 3 (Redistributive Rents).
For any $\mu\in\Delta(\Theta)$ and any strongly redistributive welfare function $w$, the two following statements are true:
  1. (i)
    If $\sigma^{\star}_{\mu}$ is feasible, then it is the unique $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation and it does not give any rent to the seller.
  2. (ii)
    If $\sigma^{\star}_{\mu}$ is not feasible, then the unique $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation gives a rent to the seller.
Proof.
A segmentation does not give any rent to the seller if and only if $p^{\star}_{\mu}$ is an optimal price in every market segment. [^31] This follows easily from the fact that the seller’s total profit is the sum of its profits on each segment. Therefore, a strongly monotone segmentation does not give a rent to the seller only if it satisfies two conditions. First, it must generate only two market segments. One containing every type weakly lower than $p^{\star}_{\mu}$, and one containing every type weakly greater than $p^{\star}_{\mu}$. Second, Ob must bind at $p^{\star}_{\mu}$ in the segment $\theta_{1}$. It can easily be seen that $\sigma^{\star}_{\mu}$ is the only strongly monotone segmentation that satisfies those two requirements. Furthermore, following Proposition˜2, for any strongly redistributive $w$, the $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is the unique strongly monotone one that is saturated. Therefore, if $\sigma^{\star}_{\mu}$ is feasible, then it is uniquely $\mathrm{P}^{\star}_{w,\mu}$-optimal and does not give a rent to the seller. On the contrary, if $\sigma^{\star}_{\mu}$ is not feasible, then the $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation must give a rent to the seller. ∎
Proposition˜3 highlights the key tradeoff generated by redistributive concerns. Benefiting poor consumers may require imposing a sacrifice on richer consumers’ surplus. Strongly redistributive welfare functions precisely characterize when the welfare gains from redistribution more than compensate for the surplus loss of higher types. Given that redistributive segmentations are efficient (Lemma˜1), as a byproduct of this loss, the seller benefits from a redistributive rent. This result is easily explainable in light of the structure of strongly redistributive segmentations (Proposition˜2). Such segmentations pool as many low-types as possible in low-price segments, allowing them to benefit from lower prices. Consequently, high types are separated from low types, enabling the seller to better price discriminate. This increases the seller’s profit, except in the special case where $\sigma^{\star}_{\mu}$ is feasible.
3.5.  Discussion
3.5.1. Redistribution and Consumer-Surplus Maximization
The results in this paper highlight a variety of ways in which redistributive segmentations can differ from consumer surplus-maximizing segmentations (bbm15). Notably, redistributive segmentations might differ from consumer surplus-maximizing (henceforth CS-maximizing) ones in how they assign consumers to segments, in the prices they generate, and in the way they distribute surplus across the buyers and the seller.
Redistributive segmentations exhibit a positive-assortative structure, assigning consumers with a lower willingness to pay to segments with low prices and consumers with a higher willingness to pay to segments with higher prices. As seen in Section˜3.3, all redistributive segmentations satisfy weak monotonicity: if a buyer of type $\theta$ consumes at price $p$, there must exist a buyer of higher type who consumes at a higher price. When the redistributive motive is sufficiently strong, the optimal redistributive segmentations also satisfy strong monotonicity: if a buyer of type $\theta$ consumes at price $p$, every buyer of a higher type must be consuming at a higher price. CS-maximizing segmentations need not satisfy even this weak notion of monotonicity. For instance, bbm15’ (bbm15) greedy algorithm may not lead to a weakly monotone segmentation (taking the equivalent direct segmentation).
The monotone structure of redistributive segmentations qualifies a sense in which redistributive segmentations are “more separating” than CS-maximizing ones. This greater separation is echoed in the distributions of prices induced by such segmentations: while CS-maximizing segmentations only induce segments with discounts relative to the uniform price, redistributive segmentations might also induce segments with prices higher than the uniform. As such, prices might be more dispersed under redistributive segmentations. Another symptom of this greater separation is the distribution of surplus across sides of the market. As treated in Section˜3.4, while CS-maximizing segmentations allocate all of the surplus generated by the segmentation to buyers, redistributive segmentations might, as a byproduct of prioritizing the welfare of poorer consumers, grant extra surplus to the seller.
Finally, it is important to note that the set of redistributive segmentations intersects with the set of CS-maximizing ones. Indeed, as a welfare function, total consumer surplus is included in the class of welfare functions we study, despite not being our focus since condition R, which captures our notion of redistributive concerns, is satisfied everywhere with equality. Regardless, from Theorem˜1 it follows that in the set of CS-maximizing segmentations, one can always find a saturated one.
The redistributive order characterized in Proposition˜1 is also helpful if one wants to use redistribution as a criterion to select among CS-maximizing segmentations. CS-maximizing segmentations cannot be improved in the redistributive order through downward transfers, but might be improvable through redistributive transfers. Departing from some CS-maximizing segmentation, sequentially performing feasible redistributive transfers will eventually result in a saturated segmentation that is also CS-maximizing.
3.5.2. Link to Stochastic Orders and Optimal Transport
Our characterization of the redistributive order, as presented in Proposition˜1, conceptually parallels the characterization of the supermodular stochastic order developed independently by Muller2013 and Meyer2015. The critical distinction lies in their respective objects of comparison: the supermodular stochastic order compares couplings—bivariate joint distributions with two prescribed marginals—whereas the redistributive order compares segmentations, which consist of joint distributions where only the first marginal (the aggregate market $\mu$) is prescribed, while the second marginal (the price distribution $\sigma_{P}$) emerges endogenously from the seller’s incentive constraints. [^32] This observation is generally valid for any persuasion problem, where the prior distribution of the state is prescribed while the distribution over the receiver’s action is determined by the obedience constraints (see Bergemann2019). As such, our characterization could be fruitfully extended to answer more generally a recurring question in the literature: When are optimal information structures monotone?
Therefore, the characterization of the supermodular order over couplings differs from that of the redistributive order over segmentations in a substantive way. The supermodular order is characterized only in terms of supermodular transfers (akin to redistributive transfers in our terminology), whereas the redistributive order is characterized in terms of both supermodular transfers, which preserve both marginals, and downward transfers, which can modify the marginal distribution of prices. This distinction arises directly from the endogenous nature of segmentations’ second marginal. Since downward transfers inevitably alter the price distribution, they cannot be applied to couplings, where both marginals are prescribed. [^33] It is noteworthy that redistributive transfers (Definition 4) correspond to the mass transfers used in optimal transport to show that the comonotone distributions are the extreme points of the transport polytope (see Carlier2022, Theorem 6.14). It could be worthwhile for future research to use similar tools to characterize the extreme points of the “persuasion polytope.”
Observe also that the optimization problem $\mathrm{P}^{\star}_{w,\mu}$ bears similarity to the discrete optimal assignment problem (see Galichon2018, Chapter 3). Again, the key difference is that rather than maximizing the expectation of a function on the set of couplings, we maximize it on the set of (efficient) segmentations.
In statistics and optimal transport theory it is known that comonotone joint distributions maximize the expectation of supermodular functions on the set of couplings (see, e.g., Theorem 2.1 in Puccetti2015, and Theorem 4.3 in Galichon2018). We show that, similarly as for couplings, the supermodularity of redistributive welfare functions induces a positive-assortative matching of willingness to pay and prices. Furthermore, we show in Lemma˜2 that saturated segmentations correspond to the maximal elements of the redistributive order, which provides a parallel to how comonotone distributions correspond to the maximal elements of the supermodular stochastic order (see claim (c) of Theorem 2.1 in Puccetti2015).
Collectively, Theorem˜1, Corollaries˜1 and 2 thus provide a natural analog for segmentations to the characterizations of (discrete) optimal assignments in optimal transport theory. However, since segmentations are constrained by the seller’s incentives, supermodularity alone no longer guarantees comonotonicity. As discussed earlier, our concept of strongly redistributive welfare functions (Definition˜6), which can be understood as a form of “strong-supermodularity,” becomes both necessary and sufficient to achieve comonotone optimal solutions.
4. Implementation Through Price Regulation
Thus far, we have analyzed the segmentation design problem from the perspective of a central planner with redistributive preferences. In practice, however, market segmentation is primarily conducted by private actors with access to consumers’ data, such as platforms. This leads to the problem of implementation: How can a regulator (she) induce the seller (he) to implement a redistributive segmentation?
If the regulator can verify the segmentation chosen by the seller, this implementation problem is trivial. The regulator can simply commit to punishing the seller if his chosen segmentation differs from her preferred one. However, it is unlikely that the regulator can perfectly observe the segmentation set by the seller. We show that observing the price distribution induced by segmentations is sufficient for the regulator to implement her preferred segmentation.
Assume that both the regulator and the seller know the aggregate market $\mu\in\Delta(\Theta)$. The regulator communicates to the seller the segmentation $\sigma$ she wishes to implement, after which the seller chooses a market segmentation $\sigma^{\prime}\in\Sigma(\mu)$. The regulator cannot observe $\sigma^{\prime}$ directly but can verify and enforce punishments based on the induced price distribution [^34] As discussed in the introduction, observing the distribution of prices is realistic; see Footnote 7. $\sigma^{\prime}_{P}$. Consequently, for any $\mu\in\Delta(\Theta)$, the set of undetectable deviations from $\sigma$ is
$\mathcal{D}_{\mu}(\sigma)\coloneqq\bigl\{\sigma^{\prime}\in\Sigma(\mu)\ |\ \sigma^{\prime}_{P}=\sigma_{P}\bigr\}.$
Since the regulator can detect any deviation that does not belong to $\mathcal{D}_{\mu}(\sigma)$, she can impose appropriate penalties and deter the seller from deviating to $\sigma^{\prime}\notin\mathcal{D}_{\mu}(\sigma)$. Hence we adopt the following notion of implementability.
Definition 8 (Price-based implementability).
Let $\mu\in\Delta(\Theta)$. A segmentation $\sigma\in\Sigma(\mu)$ is implementable through price-based regulation if there exists no $\sigma^{\prime}\in\mathcal{D}_{\mu}(\sigma)$ such that
$\sum_{p\in\operatorname{supp}(\sigma^{\prime}_{P})}p\sum_{\theta\geq p}\sigma^{\prime}(\theta,p)>\sum_{p\in\operatorname{supp}(\sigma_{P})}p\sum_{\theta\geq p}\sigma(\theta,p).$
A segmentation can be implemented through price-based regulation if there exists no strictly profitable deviation for the seller that is undetectable by the regulator. The next proposition shows that efficiency is, in fact, sufficient for implementability. [^35] Our notion of implementability is equivalent to the notion of credibility defined by linliu in the context of Bayesian persuasion. Specifically, the authors define an information structure as credible if there exists no alternative information structure that would strictly increase the sender’s payoff while unaffecting the marginal distribution of signal realizations. As such, Proposition 4 shows that any efficient segmentation is also seller-credible. Our notion of implementation is also equivalent to the notion of price-regulation defined by Schlom2024 in the context of second-degree price discrimination.
Proposition 4.
Let $\mu\in\Delta(\Theta)$. If $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$, then $\sigma$ is implementable through price-based regulation.
Proof.
Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$. Recall that, for any $(\theta,p)\in\Omega$ we must have that $\sigma(\theta,p)=\sigma(\theta\,|\,p)\,\sigma_{P}(p)$. We thus obtain the following decomposition of the platform’s profit under $\sigma$:
$\sum_{p\in\operatorname{supp}(\sigma_{P})}p\sum_{\theta\geq p}\sigma(\theta,p)$ $=\sum_{p\in\operatorname{supp}(\sigma_{P})}p\,\sigma_{P}(p)\overbrace{\sum_{\theta\geq p}\sigma(\theta\,|\,p)}^{\mathclap{\text{$=1$ since $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$}}}$
$=\sum_{p\in\operatorname{supp}(\sigma_{P})}p\,\sigma_{P}(p).$
By definition, $\sum_{\theta\geq p}\sigma^{\prime}(\theta\,|\,p)\leq\sum_{\theta\in\Theta}\sigma^{\prime}(\theta\,|\,p)=1$ for any $\sigma^{\prime}\in\Sigma(\mu)$. Therefore, for any $\sigma^{\prime}\in\mathcal{D}_{\mu}(\sigma)$, we must have in particular that
$\sum_{p\in\operatorname{supp}(\sigma^{\prime}_{P})}p\,\sigma^{\prime}_{P}(p)\sum_{\theta\geq p}\sigma^{\prime}(\theta\,|\,p)$ $\leq\sum_{p\in\operatorname{supp}(\sigma^{\prime}_{P})}p\,\sigma^{\prime}_{P}(p)$
$=\sum_{p\in\operatorname{supp}(\sigma_{P})}p\,\sigma_{P}(p)\,$
This implies that $\sigma$ is implementable through price-based regulation. ∎
Therefore, any efficient segmentation and, following Lemma˜1, any redistributive segmentation, can be implemented through price-based regulation. The proof follows from a simple observation: the difference in the seller’s profit between two segmentations that induce the same price distribution only depends on the mass of consumers that buy the good at each price. For efficient segmentations, this mass is maximal because every buyer consumes. Therefore, there cannot be any strictly beneficial deviation.
5. Concluding Remarks
Price discrimination has long been used as an instrument for redistribution. Many goods and services are provided at reduced prices for disadvantaged groups. For instance, students from low-income backgrounds often pay reduced tuition fees at universities, and it is common for the unemployed or the elderly to receive discounted prices at museums. The prevalence of such redistributive price discrimination indicates that personalized pricing is deemed socially acceptable as long as it is perceived to benefit those in need. This view is supported by experimental data on people’s attitudes to price discrimination (wu) and is expressed in policy reports (WhiteHouse2015; bourreau2018regulation). Our redistributive segmentations satisfy this social demand by ensuring that price discrimination satisfies a certain degree of price progressivity. Hence, our results inform the current policy debate over the regulation of online personalized pricing. They give qualitative guidelines regarding market segmentations that prioritize consumers with low willingness to pay.
Redistributive price discrimination is usually implemented by public and non-profit organizations (legrand; steinberg). A concern is that imposing redistributive price discrimination on the private sector could be inefficient. This paper studies how a private market with a profit-maximizing monopoly can be optimally segmented to achieve a redistributive goal. Redistributive segmentations have the twofold benefit of achieving economic efficiency (Lemma˜1) and never harming the seller—Proposition˜3 shows that redistributive segmentations can even improve the seller’s profit.
One might suggest that conditioning market segmentation on additional buyer information (e.g., income or group characteristics), rather than only the willingness to pay, would better target redistribution. Nevertheless, as highlighted in Section˜4, consumer data is usually collected and processed by private actors for which consumers’ willingness to pay is the payoff-relevant variable. It might also be impossible or illegal for regulators like competition authorities to directly access income data or protected characteristics. Hence, our optimal segmentations constitute a normative benchmark against which to evaluate welfare outcomes that could be achieved by regulating data intermediaries.
Appendices
Appendix A Proof of Proposition˜1
Proof.
For any $f,g\in\mathbb{R}^{\Omega}$ we let $\langle f,g\rangle$ be defined as
$\langle f,g\rangle=\sum_{(\theta,p)\in\Omega}f(\theta,p)\,g(\theta,p).$
We know from Lemma˜1 that we can, without loss of generality, restrict attention to segmentations supported on $\Omega$. Hence, we treat a segmentation $\sigma$ as an element of $\mathbb{R}^{\Omega}$.
[…]
$\mathcal{E}=\Bigl\{v\in\mathbb{R}^{\Omega}\ \big|\ \exists w\in\mathcal{R}\text{ and }b\in\mathbb{R}\text{ such that},\forall(\theta,p)\in\Omega,\,v(\theta,p)=w(\theta,p)+b\Bigr\}.$
[…]
Lemma 5.
[…]
Proof.
As $\mathcal{R}\subset\mathcal{E}$, if $\sigma\succsim_{\mathcal{E}}\sigma^{\prime}$, then $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$. Conversely, let $\sigma,\sigma^{\prime}$ be such that $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$, that is, for any $w\in\mathcal{R}$, $\langle\sigma,w\rangle\geq\langle\sigma^{\prime},w\rangle$. This implies that for any $w\in\mathcal{R}$ and $b\in\mathbb{R}$, $\langle\sigma,w\rangle+b\geq\langle\sigma^{\prime},w\rangle+b$. With a slight abuse of notation, by denoting $b$ also as the element of $\mathbb{R}^{\Theta\times P}$ constantly equal to $b$, $\langle\sigma,b\rangle=\langle\sigma^{\prime},b\rangle=b$. Therefore, for any $w\in\mathcal{R}$ and $b\in\mathbb{R}$,
$\langle\sigma,w\rangle+b\geq\langle\sigma^{\prime},w\rangle+b$ $\iff\langle\sigma,w\rangle+\langle\sigma,b\rangle\geq\langle\sigma^{\prime},w\rangle+\langle\sigma^{\prime},b\rangle$
$\iff\langle\sigma,w+b\rangle\geq\langle\sigma^{\prime},w+b\rangle$
This implies that $\sigma\succsim_{\mathcal{E}}\sigma^{\prime}$. ∎
[…]
Lemma 6.
For any $w\in\mathbb{R}^{\Omega}$,
$w\in\mathcal{E}\iff\forall t\in\mathcal{T},\;\langle w,t\rangle\geq 0.$
Proof.
Necessity follows from the definition of $\mathcal{R}$ and therefore of $\mathcal{E}$. Let us prove sufficiency. Let $w$ be such that $\langle w,t\rangle\geq 0$ for every $t\in\mathcal{T}$. By way of contradiction, assume that there exists $(\theta,p),(\theta,p^{\prime})\in\Omega$ such that $p<p^{\prime}$ and $w(\theta,p)<w(\theta,p^{\prime})$. Consider the downward buyer transfer $d\in\mathbb{R}^{\Omega}$ such that $d(\theta,p)=-d(\theta,p^{\prime})=1$. Then $\langle w,d\rangle<0$, a contradiction. Assume now that there exists $(\theta,p),(\theta^{\prime},p^{\prime})\in\Omega$ such that $\theta<\theta^{\prime}$, $p<p^{\prime}$, and $w(\theta,p)-w(\theta,p^{\prime})<w(\theta^{\prime},p)-w(\theta^{\prime},p^{\prime})$. Consider the redistributive buyer transfer $r\in\mathbb{R}^{\Omega}$ such that
$r(\theta,p)=r(\theta^{\prime},p^{\prime})=-r(\theta^{\prime},p)=-r(\theta,p^{\prime})=1.$
Then $\langle w,r\rangle<0$, again a contradiction. ∎
Let $\mathrm{cone}(\mathcal{T})$ denote the conic hull of $\mathcal{T}$, the smallest convex cone containing $\mathcal{T}$, equal to the set of all positive linear combinations of finitely many elements of $\mathcal{T}$ (see ok2007real, p. 430):
$\mathrm{cone}(\mathcal{T})=\left\{\sum_{t\in T}\lambda_{t}t\ \Big|\ \text{$T\subset\mathcal{T}$, $|T\rvert<\infty$ and $\lambda\in\mathbb{R}_{+}^{T}$}\right\}.$
Note that for any $t\in\mathcal{T}$ and $\lambda_{t}\geq 0$, $\lambda_{t}t\in\mathcal{T}$. Therefore,
$\mathrm{cone}(\mathcal{T})=\left\{\sum_{i=1}^{n}t_{i}\ \Big|\ \text{$n\in\mathbb{N}^{*}$ and $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$}\right\}.$
As $\mathrm{cone}(\mathcal{T})$ is closed, we can show in the next lemma that it is the dual cone of $\mathcal{E}$ (see luenberger1969optimization, p.215).
Lemma 7.
For any $\tau\in\mathbb{R}^{\Omega}$,
$\tau\in\mathrm{cone}(\mathcal{T})\iff\forall w\in\mathcal{E},\;\langle w,\tau\rangle\geq 0.$
Proof.
Assume that $\tau\in\mathrm{cone}(\mathcal{T})$. Hence, there exists $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that $\tau=\sum_{i=1}^{n}t_{i}$. Following Lemma˜6 we have that, for any $w\in\mathcal{E}$,
$\langle\tau,w\rangle=\sum_{i=1}^{n}\langle t_{i},w\rangle\geq 0.$
Assume now that $\upsilon\notin\mathrm{cone}(\mathcal{T})$. As $\mathrm{cone}(\mathcal{T})$ is closed, the Separating Hyperplane Theorem implies that there exists $w\in\mathbb{R}^{\Omega}$ such that $\langle\upsilon,w\rangle<\langle\tau,w\rangle$ for every $\tau\in\mathrm{cone}(\mathcal{T})$. Since $\mathrm{cone}(\mathcal{T})$ is a cone, it must be that $\langle\tau,w\rangle\geq 0$ for any $\tau\in\mathrm{cone}(\mathcal{T})$ as otherwise there would exist some $\alpha>0$ such that $\langle\alpha\tau,w\rangle\leq\langle\upsilon,w\rangle$. Thus, Lemma˜6 implies that $w\in\mathcal{E}$. Finally, since $0=\min_{\tau\in\mathrm{cone}(\mathcal{T})}\langle\tau,w\rangle$, we must have $\langle\upsilon,w\rangle<0$. ∎
We can now restate the theorem as follows: for any $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$, $\sigma\succsim_{\mathcal{E}}\sigma^{\prime}$ if and only if there exists $n\in\mathbb{N}^{*}$ and a sequence $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that:
$\sigma=\sigma^{\prime}+\sum_{i=1}^{n}t_{i}$ (5)
Equation˜5 is equivalent to $\sigma-\sigma^{\prime}\in\mathrm{cone}(\mathcal{T})$, and by Lemma˜7, this is equivalent to $\langle\sigma-\sigma^{\prime},w\rangle\geq 0$ for every $w\in\mathcal{E}$. ∎
Appendix B Proof of Lemma˜2
Proof.
[…]
We now prove claim (ii). For any set $A$ we let $A=\bigl\{-a\ |\ a\in A\bigr\}$. For any two sets $A$ and $B$, we let $A+B\coloneqq\bigl\{a+b\ |\ a\in A,b\in B\bigr\}$ be the Minkowski sum of $A$ and $B$, and we let AB=A+(B)A-B=A+(-B).
[…]
Given that these are closed convex sets, the separating hyperplane theorem implies that there exists $v\in\mathbb{R}^{\Omega}$ such that:
$\forall\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu),\forall t\in\mathrm{cone}(\mathcal{T}),\quad\langle\sigma^{\prime},v\rangle\leq\langle\sigma,v\rangle\leq\langle\sigma+t,v\rangle.$
Furthermore, for any $t\in\mathrm{cone}(\mathcal{T})$, $\langle\sigma,v\rangle\leq\langle\sigma+t,v\rangle\iff 0\leq\langle t,v\rangle$, which, by Lemma˜6, implies that $v\in\mathcal{E}$. Therefore, there exists $w\in\mathcal{R}$ and $b\in\mathbb{R}$ such that $v=w+b$. Hence, for any $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$,
$\langle\sigma^{\prime},v\rangle\leq\langle\sigma,v\rangle$ $\iff\langle\sigma^{\prime},w+b\rangle\leq\langle\sigma,w+b\rangle$
$\iff\langle\sigma^{\prime},w\rangle\leq\langle\sigma,w\rangle,$
which proves that $\sigma$ is $\mathrm{P}^{\star}_{w,\mu}$-optimal. ∎
Appendix C Proof of Lemma˜3
From Proposition˜1, for any $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$, there exists no $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$ such that $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$ if and only if for any $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tau\neq 0$, $\sigma+\tau\notin\Sigma^{\mathrm{eff}}(\mu)$. We first show the following equivalence.
Lemma 8.
Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$. For any $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tau\neq 0$, $\sigma+\tau\notin\Sigma^{\mathrm{eff}}(\mu)$ if and only if for any $t\in\mathcal{T}$ such that $t\neq 0$, $\sigma+t\notin\Sigma^{\mathrm{eff}}(\mu)$.
Proof.
Necessity is straightforward. Let now $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ be such that for any $t\in\mathcal{T}$ such that $t\neq 0$, $\sigma+t\notin\Sigma^{\mathrm{eff}}(\mu)$. First, this means that $\sigma$ is on the relative border of $\Sigma^{\mathrm{eff}}(\mu)$, as otherwise, for any neighborhood around $\sigma$ included in $\Sigma^{\mathrm{eff}}(\mu)$, there would exist sufficiently small transfers $t\in\mathcal{T}$ such that $\sigma+t$ is in this neighborhood. Second, this also implies that $\sigma\neq\sigma^{F}$, where $\sigma^{F}$ is the first-degree price discrimination segmentation. That is, the segmentation supported on the line $p=\theta$. This is because downward transfers could be performed on $\sigma^{F}$ without violating the obedience constraints.
We now proceed by contradiction, by assuming that there exists $\tau\in\mathrm{cone}(\mathcal{T})$, $\tau\neq 0$, such that $\sigma^{\prime}\coloneqq\sigma+\tau\in\Sigma^{\mathrm{eff}}(\mu)$, that is, $\sigma^{\prime}\succ_{\mathcal{R}}\sigma$. As in Appendix˜B, we let $A=\bigl\{-a\ |\ a\in A\bigr\}$ for any set $A$, and $A+B\coloneqq\bigl\{a+b\ |\ a\in A,b\in B\bigr\}$ and AB=A+(B)A-B=A+(-B) for any two sets $A$ and $B$. We first prove that $\Sigma^{\mathrm{eff}}(\mu)\subset\bigl(\{\sigma\}+\mathrm{cone}(\mathcal{T})\bigr)$, which is equivalent to $\bigl(\Sigma^{\mathrm{eff}}(\mu)-\{\sigma\}\bigr)\subset\mathrm{cone}(\mathcal{T})$. Given that $\sigma^{\prime}=\sigma+\tau\in\Sigma^{\mathrm{eff}}(\mu)$, for some $\tau\neq 0$ and $\tau\notin\mathcal{T}$ this means that $\bigl(\Sigma^{\mathrm{eff}}(\mu)-\{\sigma\}\bigr)\,\cap\,\operatorname{int}\bigl(\mathrm{cone}(\mathcal{T})\bigr)\neq\varnothing$. Now suppose by contradiction that $\bigl(\Sigma^{\mathrm{eff}}(\mu)-\{\sigma\}\bigr)\not\subset\mathrm{cone}(\mathcal{T})$. As $\Sigma^{\mathrm{eff}}(\mu)$ and $\mathrm{cone}(\mathcal{T})$ are closed and convex sets, there exists $x\in\Sigma^{\mathrm{eff}}(\mu)$ that is on relative border of $\mathrm{cone}(\mathcal{T})$. This means that there exists $\tilde{\sigma}\in\Sigma^{\mathrm{eff}}(\mu)$ such that $\tilde{\sigma}=\sigma+t$ for some $t\in\mathcal{T}$ (as the borders of $\mathrm{cone}(\mathcal{T})$ are elements of $\mathcal{T}$), which contradicts our premise. Hence we proved that $\Sigma^{\mathrm{eff}}(\mu)-\{\sigma\}\subset\mathrm{cone}(\mathcal{T})$; which implies that for any $\tilde{\sigma}\in\Sigma^{\mathrm{eff}}(\mu)$, there exists $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\tilde{\sigma}=\sigma+\tau$. To see that this is a contradiction, consider $\sigma^{F}$. We know that $\sigma^{F}\in\Sigma^{\mathrm{eff}}(\mu)$ and $\sigma^{F}\neq\sigma$, but $\sigma^{F}$ cannot be reached from $\sigma$ by a finite sequence of transfers in $\mathcal{T}$; that is, there exists no $\tau\in\mathrm{cone}(\mathcal{T})$ such that $\sigma^{F}=\sigma+\tau$. ∎
Equipped with Lemma˜8, we can now proceed to the proof of Lemma˜3 by proving it is equivalent that a segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is saturated and that there exists no $t\in\mathcal{T}$ such that $t\neq 0$ and $\sigma+t\in\Sigma^{\mathrm{eff}}(\mu)$. For any $\mu\in\Delta(\Theta)$, $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ and $p\in\operatorname{supp}(\sigma_{P})$, define:
$\mathcal{I}_{\sigma}(p)=\Biggl\{q\in\Theta\setminus\{p\}\;\Big|\;p\sum_{\theta\geq p}\sigma(\theta,p)=q\sum_{\theta\geq\tilde{p}}\sigma(\theta,q)\Biggr\}.$
Proof of Lemma˜3.
[…]
We now prove part (b) in Definition˜1. Let $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$. By contradiction, assume there exists $q\in\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)$ such that $q\geq p^{\prime}$ and that $q\notin\mathcal{I}_{\sigma}(p^{\prime})$. This implies that $q>p^{\prime}$, so that the greatest type in $\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)$ that is strictly lower than $q$, defined by
$\theta_{q}:=\max\Bigl\{\theta\in\Theta\ \big|\ \theta<q,\,\theta\in\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)\Bigr\},$
is well-defined. Then, for $\varepsilon>0$ sufficiently small, the segmentation that is obtained from $\sigma$ by the redistributive buyer transfer that transfers mass $\varepsilon$ from $(\theta_{q},p^{\prime})$ and $(q,p)$ to $(\theta_{q},p)$ and $(q,p^{\prime})$ satisfies Ob, a contradiction.
[…]
We now show that there exists no $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$ that can be obtained from $\sigma$ by a downward buyer transfer. By contradiction take $(\theta^{\prime},p^{\prime\prime})\in\operatorname{supp}(\sigma)$, $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p^{\prime}<p^{\prime\prime}$, $\varepsilon>0$, and suppose that $\sigma^{\prime}$ obtained from $\sigma$ by a downward buyer transfer from $(\theta^{\prime},p^{\prime\prime})$ to $(\theta^{\prime},p^{\prime})$ is in $\Sigma^{\mathrm{eff}}(\mu)$. Part (a) in Definition˜1 implies that there exists $q>p^{\prime}$ such that $q\in\mathcal{I}_{\sigma}(p^{\prime})$. If $q\leq\theta^{\prime}$, then we must have
$p^{\prime}\sum_{\theta\geq p^{\prime}}\sigma^{\prime}(\theta,p^{\prime})<q\sum_{\theta\geq q}\sigma^{\prime}(\theta,p^{\prime}),$
which contradicts $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$. Therefore, it must be that $q>\theta^{\prime}\geq p^{\prime\prime}$. Now, part (b) in Definition˜1 implies that $q\in\mathcal{I}_{\sigma}(p^{\prime\prime})$, meaning that
$p^{\prime\prime}\sum_{\theta\geq p^{\prime\prime}}\sigma^{\prime}(\theta,p^{\prime\prime})<p^{\prime\prime}\sum_{\theta\geq p^{\prime\prime}}\sigma(\theta,p^{\prime\prime})=q\sum_{\theta\geq q}\sigma(\theta,p^{\prime\prime})=q\sum_{\theta\geq q}\sigma^{\prime}(\theta,p^{\prime\prime}),$
which contradicts $\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$. ∎
Appendix D Proof of Proposition˜2
Proof.
[…]
$\theta^{\prime}=\min\bigl\{x\in\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,p)\bigr)\ |\ x>\theta_{i}\bigr\},$
$\theta_{k}=\max\operatorname{supp}\bigl(\sigma(\,\cdot\,|\,\theta_{i})\bigr).$
Let $\varepsilon>0$ be sufficiently small and $r$ be the redistributive buyer transfer such that
$r(\theta_{i},p)=r(\theta^{\prime},\theta_{i})=-r(\theta_{i},\theta_{i})=-r(\theta^{\prime},p)=\varepsilon,$
and $r=0$ otherwise. Neglecting for the moment the negative indirect impacts of $r$ by triggering a violation of the obedient constraint in the segment $\theta_{i}$, the direct impact on the welfare of types $\theta_{i}$ and $\theta^{\prime}$ involved in the transfer $r$ is
$\varepsilon\left[(w(\theta_{i},p)-w(\theta_{i},\theta_{i}))-\left(w(\theta^{\prime},p)-w(\theta^{\prime},\theta_{i})\right)\right]$ (6)
$>$ $\varepsilon\left[(w(\theta_{i},p)-w(\theta_{i},\theta_{i}))-\left(w(\theta_{i+1},p)-w(\theta_{i+1},\theta_{i})\right)\right]>0,$ (7)
[…]
$\theta\sum_{x\geq\theta}\sigma(x,\theta)<\hat{\theta}\left(\sum_{x\geq\hat{\theta}}\sigma(x,\theta)+\varepsilon\right).$
To recover the obedience of the price recommendation $\theta_{i}$, we implement the transfer $t\colon\Omega\to\mathbb{R}$, which transfers a mass $\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon$ of buyers from $(\theta_{k},\theta_{i})$ to $(\theta_{k},\theta_{k})$: that is,
$t(\theta_{k},\theta_{k})=-t(\theta_{k},\theta_{i})=\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon,$
and $t=0$ otherwise. First, for sufficiently small $\varepsilon$, $\sigma+t\in\Delta(\Omega)$. Second, implementing $t$ cannot break any obedience constraints, and it corrects from any violation of obedience constraints induced by the transfer $r$. To see this, look at the difference between the profit made by setting the price $\theta_{i}$ and setting the price $q\geq\theta_{i+1}$:
$\theta_{i}\left(\sum_{x\geq\theta_{i}}\sigma(x,\theta_{i})-\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)-q\left(\sum_{x\geq q}\sigma(x,\theta_{i})+\varepsilon-\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$
$\geq$ $\varepsilon\left(\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}(q-\theta_{i})-q\right)=\varepsilon\frac{\theta_{i}(q-\theta_{i+1})}{\theta_{i+1}-\theta_{i}}\geq 0$
where the first inequality follows from $\theta_{i}$ being an optimal price in segment $\theta_{i}$ according to $\sigma$, and the second inequality follows from $q\geq\theta_{i+1}$.
Therefore, $\sigma+r+t\in\Sigma^{\mathrm{eff}}(\mu)$. The loss from the transfer $t$ is
$\frac{\theta_{i+1}}{\theta_{i+1}-\theta}\varepsilon\bigr[w(\theta_{k},\theta_{i})-w(\theta_{k},\theta_{k})\bigr].$ (8)
Therefore, implementing the transfers $r$ and $t$ is beneficial if and only if 6 $\geq$ 8. If $w$ is strongly redistributive, we have that 7 $\geq$ 8. Given that 6 $\geq$ 7, this ends the proof.
[…]
$\bigl[w(\theta_{i},p)-w(\theta_{i},\theta_{i})\bigr]$- $\bigl[w(\theta_{i+1},p)-w(\theta_{i+1},\theta_{i})\bigr]$ (9)
$<\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\bigl[w(\theta_{k},\theta_{i})-w(\theta_{k},\theta_{k})\bigr].$
[…]
  • one segment $p$ including all types $p$ and some type $\theta_{i}$;
  • one segment $\theta_{i}$ including the remaining types $\theta_{i}$ and all types $\theta_{i+1}$;
  • and one segment $\theta_{k}$ including all types $\theta_{k}$.
Such a $\mu$ exists. For instance, let $\mu(p)>0$ and $\mu(\theta_{i+1})>0$ be arbitrarily small, and set $\mu(\theta_{i})=\frac{p}{\theta_{i}-p}\mu(p)+\frac{\theta_{i+1}-\theta_{i}}{\theta_{i}}\mu(\theta_{i+1})$ and $\mu(\theta_{k})=1-\mu(p)-\mu(\theta_{i})-\mu(\theta_{i+1})$. One can verify that this $\mu$ satisfies all the desired conditions. Let $\varepsilon>0$ be sufficiently small. We will now perform two transfers. The first one is a downward transfer $d$ defined by:
$d(\theta_{k},\theta_{i})=-d(\theta_{k},\theta_{k})=\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon,$
$d=0$ otherwise. The second transfer, $s$, is the reverse of a redistributive one. It is defined by:
$s(\theta_{i},\theta_{i})=s(\theta_{i+1},p)=-s(\theta_{i+1},\theta_{i})=-s(\theta_{i},p)=\varepsilon,$
and $s=0$ otherwise. We check that $\sigma+d+s$ satisfies Ob. First, for sufficiently small $\varepsilon$, this does not change the optimality of price $p$ in segment $p$. Second, in segment $\theta_{i}$, the transfers do not change the obedience constraint between for $\theta_{i+1}$ as
$\theta_{i}\left(\sigma(\theta_{i},\theta_{i})+\sigma(\theta_{i+1},\theta_{i})+\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$ $-\theta_{i+1}\left(\sigma(\theta_{i+1},\theta_{i})-\varepsilon+\frac{\theta_{i+1}}{\theta_{i+1}-\theta_{i}}\varepsilon\right)$
$=$ $\varepsilon\left(\frac{\theta_{i}\theta_{i+1}}{\theta_{i+1}-\theta_{1}}-\frac{\theta_{i}\theta_{i+1}}{\theta_{i+1}-\theta_{1}}\right)$
$=$ $0,$
[…]
Appendix E Proof of Lemma˜4
Proof.
First, if the uniform price of $\mu$ is $\theta_{1}$, there exists a unique segmentation that is saturated, the one that does not segment the market. Suppose now that the uniform price of $\mu$ is strictly greater than $\theta_{1}$. Let $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ be strongly monotone and saturated. Denote as $\theta_{k}$ the maximal type that is in the segment $\theta_{1}$. By Definition˜1, there exists $\theta_{i}$, with $\theta_{1}<\theta_{i}\leq\theta_{k}$ such that $\theta_{i}\in\mathcal{I}_{\sigma}(\theta_{1})$. Consequently, any segmentation that is strongly monotone and saturated must have the same market segment $\theta_{1}$. Conclude by iterating the same argument for the other segments. ∎